### v1

# KAGNNs: Kolmogorov-Arnold Networks meet Graph Learning

Abstract
In recent years, Graph Neural Networks (GNNs) have become the de facto tool for learning node and graph representations. Most GNNs typically consist of a sequence of neighborhood aggregation (a.k.a., message passing) layers. Within each of these layers, the representation of each node is updated from an aggregation and transformation of its neighbours representations at the previous layer. The upper bound for the expressive power of message passing GNNs was reached through the use of MLPs as a transformation, due to their universal approximation capabilities. However, MLPs suffer from well-known limitations, which recently motivated the introduction of Kolmogorov-Arnold Networks (KANs). KANs rely on the Kolmogorov-Arnold representation theorem, rendering them a promising alternative to MLPs. In this work, we compare the performance of KANs against that of MLPs in graph learning tasks. We perform extensive experiments on node classification, graph classification and graph regression datasets. Our preliminary results indicate that while KANs are on-par with MLPs in classification tasks, they seem to have a clear advantage in the graph regression tasks.
1 Introduction
Graphs are structural representations of information which are useful for modeling many types of data. They arise naturally in a wide range of application domains, and their abstract nature offers increased flexibility. Typically, the nodes of a graph represent entities, while the edges capture the interactions between them. For instance, in social networks, nodes represent individuals, and edges represent their social interactions. In chemo-informatics, molecules are commonly modeled as graphs, with nodes corresponding to atoms and edges to chemical bonds. In other settings, molecules can also be nodes, with edges capturing their ability to bond with one another.
In many cases where graph data is available, there exist problems that cannot be solved efficiently using conventional tools (e. g., graph algorithms) and require the use of machine learning techniques. For instance, in the field of chemo-informatics, the standard approach for estimating the quantum mechanical properties of molecules leverages computationally expensive density functional theory computations [12]. Machine learning methods could serve as a more efficient alternative to those methods. Recently, Graph Neural Networks (GNNs) have been established as the dominant approach for learning on graphs [36]. Most GNNs consist of a series of message passing layers. Within a message passing layer, each node updates its feature vector by aggregating the feature vectors of its neighbors and combining the emerging vector with its own representation.
A lot of recent work has focused on investigating the expressive power of GNNs [22]. There exist different definitions of expressive power, however, the most common definition is concerned with the number of pairs of non-isomorphic graphs that a GNN model can distinguish. Two graphs are isomorphic if there exists an edge-preserving bijection between their respective sets of nodes. In this setting, a model is more expressive than another model if the former can distinguish all pairs of non-isomorphic graphs that the latter can distinguish, along with other pairs that the latter cannot [27]. Furthermore, an equivalence has also been established between the ability of GNNs to distinguish non-isomorphic graphs and their ability to approximate permutation-invariant functions on graphs [3]. This line of work gave insights into the limitations of different models [38, 29], but also led to the development of more powerful architectures [27, 24, 26].
Most maximally-expressive GNN models rely on multi-layer perceptrons (MLPs) as their main building blocks, due to their universal approximation capabilities [5, 15]. The theorem states that any continuous function can be approximated by an MLP with at least one hidden layer, given that this layer contains enough neurons. Having said that, in practice the models suffer from several limitations due to non-convex loss functions, algorithms without convergence guarantees and a notorious lack of interpretability that hinders their applicability in several domains. Recently, Kolmogorov-Arnold Networks (KANs) [23] have emerged as promising alternatives to MLPs. They are based on the Kolmogorov-Arnold representation theorem [21] which states that a continuous multivariate function can be represented by a composition and sum of a fixed number of univariate functions. KANs substitute the learnable weights and pre-defined activation functions of MLPs, with learnable activations based on B-splines and summations. The initial results demonstrate that KANs have the potential to be more accurate than MLPs in low dimensions, while simultaneously being more interpretable.
In this paper, we present a thorough empirical comparison between GNNs that use KANs to update node representations and GNNs that utilize MLPs to that end. Our work is orthogonal to prior work that studies the expressive power of GNNs. Here, we empirically compare models that are theoretically equally expressive in terms of distinguishing non-isomorphic graphs against each other, and we study the impact of the different function approximation modules (i. e., KANs or MLPs) on the model’s performance. We evaluate the different GNN models on several standard node classification, graph classification and graph regression datasets.
The rest of this paper is organized as follows. Section 2, provides an overview of the tasks we address in this paper, as well as a description of message passing GNNs and Kolmogorov-Arnold networks. In section 3, we introduce the KAGIN (Kolmogorov-Arnold Graph Isomorphism Network) and KAGCN (Kolmogorov-Arnold Graph Convolution Network) models, which are variants of existing GNNs, and which leverage KANs to update node features within each layer. In section 4, we present extensive empirical results comparing the above models with their vanilla counterparts in several tasks. Finally, section 5 concludes the paper.
2 Background
2.1 Considered Graph Learning Tasks
Before presenting the tasks on which we focus in this study, we start by introducing some key notation for graphs. Let $\mathbb{N}$ denote the set of natural numbers, i. e., $\{1,2,\ldots\}$. Then, $[n]=\{1,\ldots,n\}\subset\mathbb{N}$ for $n\geq 1$. Let $G=(V,E)$ be an undirected graph, where $V$ is the vertex set and $E$ is the edge set. We denote by $n$ the number of vertices and by $m$ the number of edges, i. e., $n=|V|$ and $m=|E|$. Let $g\colon V\rightarrow[n]$ denote a bijective mapping from the space of nodes to set $[n]$. Let $\mathcal{N}(v)$ denote the the neighbourhood of vertex $v$, i. e., the set $\{u\mid\{v,u\}\in E\}$. The degree of a vertex $v$ is $\deg(v)=|\mathcal{N}(v)|$. Each node $v\in V$ is associated with a $d$-dimensional feature vector $\mathbf{x}_{v}\in\mathbb{R}^{d}$, and the feature matrix for all nodes is represented as $\mathbf{X}\in\mathbb{R}^{n\times d}$. Thus, $\mathbf{x}_{v}$ is equal to the $g(v)$-th row of $\mathbf{X}$.
In node classification, each node $v\in V$ is associated with a label $y_{v}$ that represents a class. The task is to learn a function that maps nodes to their class labels, i. e., to learn a function $f_{\text{node}}$ such that $f_{\text{node}}(v,G,\mathbf{X})=y_{v}$. In graph regression/classification, the dataset consists of a collection of $N$ graphs $G_{1},\ldots,G_{N}$ along with their class labels/targets $y_{G_{1}},\ldots,y_{G_{N}}$. The task is then to learn a function that maps graphs to their class labels/targets, i. e., to learn a function $f_{\text{graph}}$ such that $f_{\text{graph}}(G,\mathbf{X})=y_{G}$, which can be discrete or continuous, for graph classification or graph regression, respectively.
The standard approach for learning such predictors (both for node- and graph-level tasks) is to first embed the nodes of the graph(s) into some vector space. That is, we aim to learn $\mathbf{H}=\texttt{EMBEDDING}(G,\textbf{X})\in\mathbb{R}^{n\times d_{e}}$ where $d_{e}$ denotes the embedding dimension. Then, the $g(v)$-th row of matrix $\mathbf{H}$ represents the embedding of node $v$. Let $\mathbf{h}_{v}$ denote this embedding. For node-level tasks, we can use $\mathbf{h}_{v}$ to predict directly the class label/target of node $v$. For graph-level tasks, we also need to apply a readout function on all the representations of the graph’s nodes to obtain a representation $\mathbf{h}_{G}=\texttt{READOUT}(\mathbf{H})$ for the entire graph.
One particularly desirable property of such models is permutation invariance. That is, the embedding $\mathbf{h}_{G}$ of a graph needs to be the same regardless of the ordering of its nodes . Indeed, these orderings do not hold any semantic meaning and different orderings give rise to isomorphic graphs. Permutation invariance is achieved at the readout step by utilizing a permutation invariant operation over the rows of $\mathbf{H}$, such as the sum, max or mean operators.
2.2 Graph Neural Networks
One of the most widely-used paradigms for designing such permutation invariant models is the message passing framework [12] which consists of a sequence of layers and whithin each layer the embedding of each node is computed as a learnable function of its neighbors’ embeddings. Formally, the embedding $\mathbf{h}_{v}^{(\ell)}\in\mathbb{R}^{d_{\ell}}$ at layer $\ell$ is computed as follows:
$\mathbf{h}_{v}^{(\ell)}=\phi^{(l)}\left(\mathbf{h}_{v}^{(\ell-1)},\bigoplus\limits_{u\in\mathcal{N}(v)}\mathbf{h}_{u}^{(\ell-1)}\right)$ (1)
where $\bigoplus$ is a permutation-invariant aggregation function (e. g., mean, sum), and $\phi^{(\ell)}$ is a differentiable function (e. g., linear transformation, MLP) that combines and transforms the node’s previous embedding with the aggregated vector of its neighbors.
As discussed above, in this paper, we focus on the functions that different GNN models employ to update node representations. Many existing GNNs use a $1$-layer perceptron (i. e., a linear mapping followed by a non-linear activation function) within each neighborhood aggregation layer to update node features [7, 20, 40]. For instance, each layer of the the Graph Convolutional Network (GCN) [20] is defined as follows:
$\mathbf{h}_{v}^{(\ell)}=\sigma\left(\mathbf{W}^{(\ell)}\sum\limits_{u\in\mathcal{N}(v)\cup\{v\}}\frac{\mathbf{h}_{u}^{(\ell-1)}}{\sqrt{(\deg(v)+1)(\deg(u)+1)}}\right)$ (2)
where $\sigma$ is a non-linear activation and $\mathbf{W}^{(\ell)}$ is a trainable weight matrix.
However, the 1-layer perceptron is not a universal approximator of multiset functions [38]. Thus, the emerging GNN might not be expressive enough for some tasks. Thus, more recent models use MLPs instead of 1-layer perceptrons to update node representations [2, 6, 28, 29]. It is well-known that standard message passing GNNs are bounded in expressiveness by the Weisfeiler-Leman (WL) test of isomorphism [38]. While two isomorphic graphs will always be mapped to the same representation by such a GNN, some non-isomorphic graphs might also be assigned identical representations.
A model that can achieve the same expressive power as the WL test, given sufficient width and depth of the MLP, is the Graph Isomorphism Network (GIN) [38], which is defined as follows:
$\mathbf{h}^{(\ell)}_{v}=\texttt{MLP}^{(\ell)}\left((1+\epsilon^{(\ell)})\cdot\mathbf{h}^{(\ell-1)}_{v}+\sum_{u\in\mathcal{N}(v)}\mathbf{h}^{(\ell-1)}_{u}\right)$ (3)
where $\epsilon^{(\ell)}$ denotes a trainable parameter, and $\texttt{MLP}^{(\ell)}$ a trainable MLP.
The GIN model can achieve its full potential if proper weights (i. e., for the different $\texttt{MLP}^{(\ell)}$ layers and $\epsilon^{(\ell)}$) are learned. However, in practice, GIN might fail to learn those weights due to limited training data and due to limitations of the employed training algorithm (e. g., stochastic gradient descent). This has motivated a series of works which focused on improving the training procedure of GNNs. For example, Ortho-GConv is an orthogonal feature transformation that can address GNNs’ unstable training [13]. Other works have studied how to initialize the weights of the MLPs of the message passing layers of GNNs. It was shown that by adopting the weights of converged MLPs as the weights of corresponding GNNs can lead to performance improvements in node classification tasks [14]. On the other hand, there exist settings where there is no need of complex learning models. This has led to the development of methods for simplifying GNNs. This can be achieved by removing the nonlinearities between the neighborhood aggregation layers and collapsing the resulting function into a single linear transformation [35] or by feeding the node features into a neural network which generates predictions and then propagate those predictions via a personalized PageRank scheme [9].
2.3 Kolmogorov-Arnold Networks
Presented as an alternative to the MLP, the Kolmogorov-Arnold Network (KAN) architecture has recently attracted a lot of attention in the machine learning community [23]. As mentioned above, this model relies on the Kolmogorov-Arnold representation theorem, which states that any multivariate function $f:\left[0,1\right]^{d}\rightarrow\mathbb{R}$ can be written as:
$f(\mathbf{x})=\sum_{i=1}^{2d+1}\Phi_{i}\left(\sum_{j=1}^{d}\phi_{ij}(\mathbf{x}_{j})\right)$ (4)
where all $\Phi_{\square}$ and $\phi_{\square}$ functions are univariate, and the sum is the only multivariate operator.
Equation (4) can be seen as a two-step process. First, a different set of univariate non-linear activation functions is applied to each dimension of the input, and then the output of those functions are summed up. The authors rely on this interpretation to define a Kolmogorov-Arnold Network (KAN) layer, which is a mapping between a space $A\subseteq\mathbb{R}^{d}$ and a different space $B\subseteq\mathbb{R}^{d^{\prime}}$, (identical in use to an MLP layer). Such a layer consists of $d\times d^{\prime}$ trainable functions $\{\phi_{ij},~{}1\leq i\leq d^{\prime},~{}1\leq j\leq d\}$. Then, for $\mathbf{x}\in A$, we compute its image $\mathbf{x}^{\prime}$ as:
$\mathbf{x}_{i}^{\prime}=\sum\limits_{j=1}^{d}\phi_{ij}(\mathbf{x}_{j})$ (5)
Stacking two such layers, one with input dimension $d$ and output dimension $12d+$ 1, and another with input dimension $12d+$ 1 and output dimension $1$, we obtain Equation (4), and the derived model is a universal function approximator.
This seemingly offers a complexity advantage compared to MLPs, since the number of univariate functions required to represent any multivariate function from $\left[0,1\right]^{d}$ to $\mathbb{R}^{d^{\prime}}$ is at most $(2d^{2}+d)\times d^{\prime}$, whereas the universal approximation theorem for the MLP requires a possibly infinite number of neurons. However, as stated by the original paper, the behavior of such univariate functions might be arbitrarily complex (e. g., fractal, non-smooth), thus leading to them being non-representable, and non-learnable.
MLPs relax the arbitrary-width constraint by stacking finite-width layers. Likewise, KANs relax the arbitrary-complexity constraints on the non-linearities by stacking KAN layers. Thus, the output of a function is given by:
$y=\texttt{KAN}(\mathbf{x})=\Phi_{L}\circ\Phi_{L-1}\circ\cdots\circ\Phi_{1}(\mathbf{x})$ (6)
where $\Phi_{1},\ldots,\Phi_{L}$ are KAN layers.
The original paper uses splines (i. e., trainable piecewise-polynomial functions) as nonlinearities. This allows to retain a high expressivity for a relatively small number of parameters, at the cost of enforcing some local smoothness. A layer $\ell$ is thus a $d_{\ell}\times d_{\ell-1}$ grid of splines. The degree used for each spline (called spline order), as well as the number of splines used for each function (called grid size) are both hyperparameters of the architecture.
Even though KANs were introduced very recently, they have already been applied to different problems such as in the task of satellite image classification [4] and for predicting the pressure and flow rate of flexible electrohydrodynamic pumps [31]. So far, most efforts have focused on time series data [39]. For instance, KANs have been evaluated in the satellite traffic forecasting task [34]. Furthermore, they have been combined with architectures that are traditionally leveraged in time series forecasting tasks such as the Long Short-Term Memory Network [11] and the Transformer [10]. The work closest to ours is the one reported in [37], where the authors propose FourierKAN-GCF. This is a GNN model designed for the task of graph collaborative filtering where the feature transformation in the neighborhood aggregation layers is performed by KANs.
3 KAN-based GNN Layers
We next derive variants of the GIN and GCN models which use KANs to transform the node features instead of fully-connected layers or MLPs.
3.1 The KAGIN Layer
To achieve its maximal expressivity, the GIN model relies on the MLP architecture and its universal approximator property. Since KAN is also a universal function approximator, we could achieve the same expressive power using KANs in lieu of MLPs. We thus propose the KAGIN model which is defined as follows:
$\mathbf{h}^{(\ell)}_{v}=\texttt{KAN}^{(l)}\left((1+\epsilon)\cdot\mathbf{h}^{(\ell-1)}_{v}+\sum_{u\in\mathcal{N}(v)}\mathbf{h}^{(\ell-1)}_{u}\right)$ (7)
With theoretically-sound KANs (i. e., with arbitrarily complex components), this architecture is exactly as expressive as the vanilla GIN model with arbitrary layer width. While this is not guaranteed with the spline-based implementation with limited grid size, the empirical results in the original paper demonstrate the great expressive power of KANs [23], especially of small models and setting where regularity is expected.
3.2 The KAGCN Layer
GCN-based architectures have achieved great success in node classification tasks. While in our experiments we evaluate KAGIN on node classification datasets, the objective advantage of GCN over GIN on some them does not facilitate a fair estimation of KANs’ potential in this context. To this end, we also propose a variant of the GCN model. Specifically, we substitute the parameters and ReLU function of the standard GCN [20] model with a single KAN layer (defined in Equation (6)) to obtain the KAGCN layer:
$\mathbf{h}^{(\ell)}_{v}=\Phi^{(\ell)}\left(\sum\limits_{u\in\mathcal{N}(v)\cup\{v\}}\frac{\mathbf{h}_{u}^{(\ell-1)}}{\sqrt{(\deg(v)+1)(\deg(u)+1)}}\right)$ (8)
where $\Phi^{(\ell)}$ denotes a single KAN layer. In the familiar matrix formulation, where $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ is the adjacency matrix with self-loops and $\tilde{\mathbf{D}}$ the diagonal degree matrix of $\tilde{\mathbf{A}}$, the node update rule of KAGCN can be written as:
$\mathbf{H}^{(\ell)}=\Phi^{(\ell)}\Big{(}\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}\mathbf{H}^{(\ell-1)}\Big{)}$ (9)
where the different rows of $\mathbf{H}^{(\ell)}$ store the representations of the different nodes of the graph.
4 Empirical Evaluation
In this section, we compare the KAGIN and KAGCN models with the GIN and GCN models in the following tasks: node classification, graph classification and graph regression. The code for reproducing the results is available at https://github.com/RomanBresson/KAGNN. All models are implemented with PyTorch [30]. For KAN layers, we rely on a publicly available implementation [^1] https://github.com/Blealtan/efficient-kan.
4.1 Node classification
Datasets.
To evaluate the performance of GNNs with KAN layers in the context of node classification, we use $7$ well-known datasets of varying sizes and types, including homophilic (Cora, Citeseer [20] and Ogbn-arxiv [16]) and heterophilic (Cornell, Texas, Wisconsin, Actor) networks. The homophilic networks are already split into training, validation and test sets, while the heterophilic datasets are accompanied by fixed $10$-fold cross validation indices.
Experimental setup.
For every dataset and model, we tune the values of the hyperparameters. Specifically, we choose the values that perform best on the validation set (i. e., lowest validation error). To find these values, we use the Optuna package [1]. We set the number of iterations of Optuna equal to $100$ trials.
For all models, the learning rate is chosen from $[10^{-3},10^{-2}]$, the number of message passing layers from $\{1,2,3,4\}$, the hidden dimension size from $\{8,9,\ldots,128\}$ and the weight decay from $\{0,0.0005\}$. For KAGIN and KAGCN, we also choose the grid size from $\{3,4,5\}$ and the spline order from $\{1,2,3,4\}$. Once the best hyperparameter values are found, we evaluate the models on the test set. For the homophilic networks, we initialize and train $10$ different models (we use $10$ different random seeds). We then evaluate the $10$ models on the test set and report the average accuracy. For the heterophilic datasets, we tune each model’s hyperparameters within each fold and we report the average accuracy across the $10$ folds.
Table 1: Average classification accuracy ($\pm$ standard deviation) of the KAGIN, GIN, KAGCN, and GCN models on the $7$ node classification datasets.
Dataset KAGIN GIN KAGCN GCN
Cora 0.7620 $\pm$ 0.0077 0.6003 $\pm$ 0.1100
0.7826
$\pm$ 0.0177
0.7447 $\pm$ 0.0638
Citeseer
0.6837
$\pm$ 0.0117
0.4748 $\pm$ 0.0574 0.6409 $\pm$ 0.0185 0.6680 $\pm$ 0.0110
Ogbn-arxiv 0.2450 $\pm$ 0.0806 0.2396 $\pm$ 0.0718 0.2997 $\pm$ 0.0274
0.5346
$\pm$ 0.0581
Cornell
0.5243
$\pm$ 0.0813
0.4270 $\pm$ 0.1171 0.4892 $\pm$ 0.0547 0.4568 $\pm$ 0.0667
Texas 0.5973 $\pm$ 0.0505
0.6054
$\pm$ 0.0516
0.5811 $\pm$ 0.0440 0.5838 $\pm$ 0.0386
Wisconsin 0.5510 $\pm$ 0.0720 0.5216 $\pm$ 0.1336 0.5196 $\pm$ 0.1020
0.5765
$\pm$ 0.0828
Actor 0.2890 $\pm$ 0.0139 0.2820 $\pm$ 0.0108 0.2755 $\pm$ 0.0134
0.3214
$\pm$ 0.0167
Results.
The results are given in Table 1. We can see that KAGIN outperforms GIN on all but one datasets. On some datasets, the difference in performance between the two models is significant. For example, on Citeseer and Cora, KAGIN offers a respective absolute improvement of $20.89\%$ and $16.17\%$ in accuracy over GIN. On the other hand, KAGCN outperforms GCN on only $2$ out of the $7$ considered datasets. On some datasets (e. g., Ogbn-arxiv), GCN significantly outperforms KAGCN. With regards to the different families of models, there is no clear winner. The KAGIN and GIN models are the best performing methods on $3$ datasets, while the KAGCN and GCN models achieve the highest accuracy on the rest of the datasets. Overall, the results in the node classification experiments are mixed.
The use of KANs in the GIN architecture brings performance improvements on $6/7$ datasets, but $5/7$ experiments overlap in the confidence intervals. Such an overlap also occurs in $4$ experiments of the GCN-based models, but overall we can contend that KAN has a more positive impact on the GIN architecture than on the GCN architecture.
With regards to the optimal hyperparameters, we found that KAGIN required a larger grid size on average compared to KAGCN. In general, a grid size of $4$ was chosen in $8/14$ experiments and the predominant spline order was $1$, appearing in $8/14$ experiments while $3$ appears $2$ times and $4$ only once. The number of hidden layers and their sizes varied significantly through datasets but overall KAGIN had a substantially larger average hidden layer size i. e., $75.4$ compared to $48$ for KAGIN, as $128$ was chosen for $3$ datasets compared to only $1$ for KAGCN. This is in contrast to previous findings contending that GIN-based models requires more complex learning procedures, a pattern that seems to withstand with KANs.
Training times.
We present in Table 2 the training time per epoch for different configurations of the KAGIN and GIN models. We observe that for a given number of message passing layers and hidden dimension size, the KAGIN model is computationally more expensive than GIN. If the grid size and spline order hyperparameters of KAGIN are set to $1$, the difference in running time between the two models is very small (less than $0.02$ seconds per epoch for all configurations). However, the complexity of KAGIN increases are the grid size and spline order increase. Overall, our results suggest that the running time of KAGIN is slightly greater than that of GIN, and by no means prohibitive.
Table 2: Training time per epoch on the Ogbn-arxiv dataset.
Architecture
Message Passing
Layers
Hidden
Dimension
Grid Size Spline Order # Parameters
Training time
(s/epoch)
GIN 2 16 NA NA 5336 0.035
GIN 2 64 NA NA 27368 0.041
GIN 4 16 NA NA 6936 0.037
GIN 4 64 NA NA 52200 0.056
KAGIN 2 16 1 1 31232 0.046
KAGIN 2 64 1 1 63488 0.048
KAGIN 2 16 4 1 54656 0.059
KAGIN 2 16 4 3 70272 0.146
KAGIN 4 16 1 1 38400 0.056
KAGIN 4 64 1 1 116736 0.072
KAGIN 4 64 4 1 204288 0.109
KAGIN 4 64 4 3 262656 0.298
4.2 Graph Classification
Datasets.
In this set of experiments, we compare the KAGIN model against GIN on standard graph classification benchmark datasets [25]. We experiment with the 7 following datasets: (1) MUTAG, (2) DD, (3) NCI1, (4) PROTEINS, (5) ENZYMES, (6) IMDB-B, (7) IMDB-M. The first $5$ datasets come from bio- and chemo-informatics, while the last $2$ are social interaction datasets.
Experimental setup. We follow the experimental protocol proposed in [8]. Thus, we perform $10$-fold cross-validation to obtain an estimate of the generalization performance of each method, while within each fold a model is selected based on a $90\%/10\%$ split of the training set. We use the splits provided in [8]. We use the Optuna package to select the model that achieves the lowest validation error. We set the number of iterations of Optuna equal to $100$.
For a fair comparison, we set the number of message passing layers of both models to a fixed value for each dataset. Based on preliminary experiments, on MUTAG, PROTEINS, IMDB-B and IMDB-M, we set the number of layers to $2$. On DD, we set it to $3$. On ENZYMES, we set it to $4$ and finally, on NCI1, we set it to $5$. To produce graph representations, we use the sum operator. The produced graph representations are fed to an MLP (for GIN) or a KAN (for KAGIN) layer which computes the output. We train each model for $1,000$ epochs by minimizing the cross entropy loss. We use the Adam optimizer for model training [19]. We apply batch normalization [17] and dropout [33] to the output of each message passing layer. We also use early stopping with a patience of $20$ epochs. For both models, we choose the number of hidden layers from $\{2,3,4\}$ and the dropout rate from $[0.0,0.5]$. For GIN, we choose the hidden dimension size from $\{8,9,\ldots,256\}$ and the learning rate from $[10^{-5},10^{-2}]$. For KAGIN, we choose the hidden dimension size from $\{2,3,\ldots,128\}$, the grid size from $\{1,2,\ldots,16\}$, the spline order from $\{1,2,\ldots,8\}$ and the learning rate from $[10^{-4},10^{-2}]$. For the IMDB-B and IMDB-M datasets, where there are no node features, we annotate nodes with one-hot encodings of their degrees, up to $35$ (all degrees above $35$ are set equal to $35$). Once the best hyperparemeters are found for a given split, we train $3$ different models on the training set of the split (in order to limit the impact of random initialization) and evaluate them on the test set of the split. This yields $3$ test accuracies, and we compute their average.
Table 3: Average classification accuracy ($\pm$ standard deviation) of the KAGIN and GIN models on the $6$ graph classification datasets.
MUTAG DD NCI1 PROTEINS
GIN 85.09 $\pm$ 5.97 73.68 $\pm$ 3.89 79.57 $\pm$ 1.27 70.26 $\pm$ 3.62
KAGIN
85.45
$\pm$ 7.38
75.46
$\pm$ 4.16
79.72
$\pm$ 1.79
71.78
$\pm$ 3.12
ENZYMES IMDB-B IMDB-M
GIN
53.72
$\pm$ 9.23
73.03 $\pm$ 2.86
49.78
$\pm$ 3.38
KAGIN 42.94 $\pm$ 5.69
73.83
$\pm$ 4.35
49.78
$\pm$ 4.69
Results.
Table 3 illustrates the average classification accuracies and the corresponding standard deviations of the two models on the different datasets. We observe that the two models achieve similar levels of performance. KAGIN outperforms GIN on $5$ out of the $7$ datasets. However, the difference in performance between the two models is very small. This suggests that the two models are similar in terms of expressive power. On ENZYMES, however, KAGIN was found to perform much worse than GIN. Note that ENZYMES consists of more classes ($6$ classes in total) than the rest of the datasets, while ENZYMES and PROTEINS are the only datasets where the nodes of the graphs are annotated with continuous features (on the rest of the datasets, they are annotated with one-hot encodings). We hypothesize that the difference in performance is due to the inability of KANs to handle those continuous features. We thus normalized the node features within each fold by removing the mean of each feature (computed from the training samples) and then dividing by the corresponding standard deviation. We re-conducted the experiment and the average accuracy increased by approximately $6\%$ ($48.77\%$ instead of $42.94\%$). Therefore, it turns out that in some settings it might be harder for KAN layers to handle continuous features than for MLPs.
Training times.
We give in Table 4 an overview of the training times for different GIN and KAGIN architectures. We provide the number of parameters of each architecture. We notice that, for the same number of parameters, KAN is slower than its MLP counterpart. This is particularly sensitive to grid size and spline order. This makes sense since, using splines, each parameter involves more complex computations than the usual multiplication/summation of traditional MLP neurons. Moreover, some of this performance difference might come from how optimized the implementation is. It would be an interesting next step to study the relation of size to performance for KANs and MLPs, since, intuitively, the expressivity of splines should allow for smaller networks.
Table 4: Training time per epoch on NCI1 dataset (i. e., only the forward and backward pass, looped over all batches, averaged over $50$ epochs). All compared models consist of $5$ message passing layers, and of $2$ hidden layers in their respective MLP/KAN. Batch size is set to $128$.
Architecture
Hidden
Dimension
Grid Size Spline Order # Parameters
Training time
(s/epoch)
GIN 16 NA NA 3522 0.368
GIN 64 NA NA 44802 0.372
GIN 256 NA NA 670722 0.433
GIN 512 NA NA 2652162 0.742
KAGIN 16 1 1 6752 0.474
KAGIN 32 1 1 21696 0.475
KAGIN 64 1 1 76160 0.475
KAGIN 256 1 1 1091072 0.732
KAGIN 16 1 3 10048 0.658
KAGIN 32 1 3 32384 0.659
KAGIN 64 1 3 113920 0.792
KAGIN 256 1 3 1635328 2.159
KAGIN 16 4 1 11696 0.476
KAGIN 32 4 1 37728 0.478
KAGIN 64 4 1 132800 0.485
KAGIN 256 4 1 1907456 1.074
KAGIN 16 4 3 14992 0.662
KAGIN 32 4 3 48416 0.726
KAGIN 64 4 3 170560 1.005
KAGIN 256 4 3 2451712 3.092
4.3 Graph Regression
Datasets.
We experiment with two molecular datasets: (1) ZINC-12K [18], and (2) QM9 [32]. ZINC-12K consists of $12,000$ molecules. The task is to predict the constrained solubility of molecules, an important chemical property for designing generative GNNs for molecules. The dataset is already split into training, validation and test sets ($10,000$, $1,000$ and $1,000$ graphs in the training, validation and test sets, respectively). QM9 contains approximately $134,000$ organic molecules. Each molecule consists of Hydrogen (H), Carbon (C), Oxygen (O), Nitrogen (N), and Flourine (F) atoms and contain up to 9 heavy (non Hydrogen) atoms. The task is to predict $12$ target properties for each molecule. The dataset was divided into a training, a validation and a test set according to a $80\%/10\%/10\%$ split.
Experimental setup.
We perform grid search to select values for the different hyperparameters. For both models, we choose the number of hidden layers from $\{2,3,4\}$, and the learning rate from $\{10^{-3},10^{-4}\}$. For GIN, we choose the hidden dimension size from $\{32,64,128,256,512,1024\}$, while for KAGIN, we choose it from $\{4,8,16,32,64,128,256\}$. For KAGIN, we also select the grid size from $\{1,3,5,8,10\}$ and the spline order from $\{3,5\}$. To produce graph representations, we use the sum operator. The emerging graph representations are finally fed to an MLP (for GIN) or a KAN (for KAGIN) layer which produces the output. We train each model for $1,000$ epochs by minimizing the mean absolute error (MAE). We use the Adam optimizer for model training [19]. We also use early stopping with a patience of $20$ epochs. For ZINC-12K, we also use an embedding layer which maps node features into $100$-dimensional vectors. We choose the configuration that achieves the lowest validation error. Once the best configuration is found, we run $10$ experiments and report the average performance on the test set. For both datasets and models, we set the number of message passing layers to $4$. On QM9, we performed a joint regression of the $12$ targets.
Table 5: Average MAE ($\pm$ standard deviation) of the KAGIN and GIN models on the ZINC-12K and QM9 datasets.
ZINC-12K QM9
GIN 0.4131 $\pm$ 0.0215 0.0969 $\pm$ 0.0017
KAGIN
0.3000
$\pm$ 0.0332
0.0618
$\pm$ 0.0007
Results.
The results are shown in Table 5. We observe that on both considered datasets, KAGIN significantly outperforms the GIN model. Note that these datasets are significantly larger (in terms of number of samples) compared to the graph classification datasets of Table 3. More specifically, KAGIN offers an absolute improvement of approximately $0.11$ and $0.03$ in MAE over GIN. Those improvements suggest that KANs might be more effective than MLPs in regression tasks.
5 Conclusion
In this paper, we have investigated the potential of Kolmogorov-Arnold networks in graph learning tasks. Since the KAN architecture is a natural alternative to the MLP, we developed two GNN architectures, KAGCN and KAGIN, respectively analogous to the GCN and GIN models. We then compared those architectures against each other in both node- and graph-level tasks. In the classification tasks, there does not appear to be a clear winner, with both architectures outperforming each other on some datasets. In the graph regression task, however, preliminary results seem to indicate that KAN has an advantage over MLP. This paper shows, through its preliminary results, that such KAN-based GNNs are valid alternatives to the traditional MLP-based models. We thus believe that these models deserve the attention of the graph machine-learning community.
Finally, we discuss potential advantages that KANs might have over MLPs, and leave their investigation for future work. First, their ability to accurately fit smooth functions could prove highly relevant on datasets where variables interact with some regular patterns. Second, their interpretability could be leveraged to provide explanations on learned models, giving insights into the nature of interactions among entities. Finally, a thorough study of the effect of the different hyperparameters could be leveraged, allowing to fully exploit the richness of splines while retaining small networks.
6 Acknowledgements
This work was partially supported by the The Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation and partner Swedish universities and industry.

### v2

@@
 Our preliminary results indicate that while KANs are on-par with MLPs in classification tasks, they seem to have a clear advantage in the graph regression tasks.
+Code is available at https://github.com/RomanBresson/KAGNN.
 1 Introduction
@@
 GCN-based architectures have achieved great success in node classification tasks.
-While in our experiments we evaluate KAGIN on node classification datasets, the objective advantage of GCN over GIN on some them does not facilitate a fair estimation of KANs’ potential in this context.
+While in our experiments we evaluate KAGIN on node classification datasets, the objective advantage of GCN over GIN on some of the datasets does not facilitate a fair estimation of KANs’ potential in this context.
 To this end, we also propose a variant of the GCN model.
@@
 Dataset KAGIN GIN KAGCN GCN
-Cora 0.7620 $\pm$ 0.0077 0.6003 $\pm$ 0.1100
+Cora 0.7620 $\pm$ 0.0077 0.5528 $\pm$ 0.1112
 0.7826
 $\pm$ 0.0177
-0.7447 $\pm$ 0.0638
-Citeseer
-0.6837
-$\pm$ 0.0117
-0.4748 $\pm$ 0.0574 0.6409 $\pm$ 0.0185 0.6680 $\pm$ 0.0110
-Ogbn-arxiv 0.2450 $\pm$ 0.0806 0.2396 $\pm$ 0.0718 0.2997 $\pm$ 0.0274
-0.5346
-$\pm$ 0.0581
+0.7566 $\pm$ 0.0718
+Citeseer 0.6837 $\pm$ 0.0117 0.4748 $\pm$ 0.0574 0.6409 $\pm$ 0.0185
+0.6891
+$\pm$ 0.0074
+Ogbn-arxiv 0.2450 $\pm$ 0.0806 0.2110 $\pm$ 0.0628
+0.2997
+$\pm$ 0.0274
+0.2777$\pm$ 0.0699
 Cornell
@@
 $\pm$ 0.0813
-0.4270 $\pm$ 0.1171 0.4892 $\pm$ 0.0547 0.4568 $\pm$ 0.0667
+$0$$\pm$ 0.1414 0.4892 $\pm$ 0.0547 0.4162 $\pm$ 0.0737
 Texas 0.5973 $\pm$ 0.0505
@@
 $\pm$ 0.0516
-0.5811 $\pm$ 0.0440 0.5838 $\pm$ 0.0386
-Wisconsin 0.5510 $\pm$ 0.0720 0.5216 $\pm$ 0.1336 0.5196 $\pm$ 0.1020
-0.5765
-$\pm$ 0.0828
-Actor 0.2890 $\pm$ 0.0139 0.2820 $\pm$ 0.0108 0.2755 $\pm$ 0.0134
-0.3214
-$\pm$ 0.0167
+0.5811 $\pm$ 0.0440 0.5892 $\pm$ 0.0415
+Wisconsin
+0.5510
+$\pm$ 0.0720
+0.4961 $\pm$ 0.1532 0.5196 $\pm$ 0.1020 0.5216 $\pm$ 0.0729
+Actor
+0.2890
+$\pm$ 0.0139
+0.2817 $\pm$ 0.0110 0.2755 $\pm$ 0.0134 0.2852 $\pm$ 0.0129
 Results.
@@
 On some datasets, the difference in performance between the two models is significant.
-For example, on Citeseer and Cora, KAGIN offers a respective absolute improvement of $20.89\%$ and $16.17\%$ in accuracy over GIN.
-On the other hand, KAGCN outperforms GCN on only $2$ out of the $7$ considered datasets.
-On some datasets (e.
-g., Ogbn-arxiv), GCN significantly outperforms KAGCN.
-With regards to the different families of models, there is no clear winner.
-The KAGIN and GIN models are the best performing methods on $3$ datasets, while the KAGCN and GCN models achieve the highest accuracy on the rest of the datasets.
-Overall, the results in the node classification experiments are mixed.
-The use of KANs in the GIN architecture brings performance improvements on $6/7$ datasets, but $5/7$ experiments overlap in the confidence intervals.
-Such an overlap also occurs in $4$ experiments of the GCN-based models, but overall we can contend that KAN has a more positive impact on the GIN architecture than on the GCN architecture.
+For example, on Citeseer and Cora, KAGIN offers a respective absolute improvement of $20.89\%$ and $20.9\%$ in accuracy over GIN.
+On the other hand, KAGCN outperforms GCN on $3/7$ considered datasets, albeit being the best in $2/3$ homophilic networks and exhibiting very similar accuracy in $3/4$ heterophilic ones.
+The KAGIN and GIN models exhibit the best performance in heterophilic networks while the KAGCN and GCN models achieve the highest accuracy on the homophilic.
+The use of KANs in the GIN architecture brings performance improvements on $6/7$ datasets, but $5/7$ experiments overlap in the confidence intervals, an overlap that is more prevalent with GCN-based models in $6/7$.
+Overall we can contend that KAN has a more positive impact on the GIN architecture than on the GCN architecture.
+Moreover, the introduction of KAN does not alleviate the inherent disadvantage of the model e.g.
+GCN accuracy with heterophilic networks [41], however, it improves the models in the majority of the results i.e.
+$5/7$.
 With regards to the optimal hyperparameters, we found that KAGIN required a larger grid size on average compared to KAGCN.
@@
 (s/epoch)
-GIN 2 16 NA NA 5336 0.035
-GIN 2 64 NA NA 27368 0.041
-GIN 4 16 NA NA 6936 0.037
-GIN 4 64 NA NA 52200 0.056
-KAGIN 2 16 1 1 31232 0.046
-KAGIN 2 64 1 1 63488 0.048
-KAGIN 2 16 4 1 54656 0.059
-KAGIN 2 16 4 3 70272 0.146
-KAGIN 4 16 1 1 38400 0.056
-KAGIN 4 64 1 1 116736 0.072
-KAGIN 4 64 4 1 204288 0.109
-KAGIN 4 64 4 3 262656 0.298
+GIN 2 16 NA NA 5,336 0.035
+GIN 2 64 NA NA 27,368 0.041
+GIN 4 16 NA NA 6,936 0.037
+GIN 4 64 NA NA 52,200 0.056
+KAGIN 2 16 1 1 31,232 0.046
+KAGIN 2 64 1 1 63,488 0.048
+KAGIN 2 16 4 1 54,656 0.059
+KAGIN 2 16 4 3 70,272 0.146
+KAGIN 4 16 1 1 38,400 0.056
+KAGIN 4 64 1 1 116,736 0.072
+KAGIN 4 64 4 1 204,288 0.109
+KAGIN 4 64 4 3 262,656 0.298
 4.2 Graph Classification
@@
 (s/epoch)
-GIN 16 NA NA 3522 0.368
-GIN 64 NA NA 44802 0.372
-GIN 256 NA NA 670722 0.433
-GIN 512 NA NA 2652162 0.742
-KAGIN 16 1 1 6752 0.474
-KAGIN 32 1 1 21696 0.475
-KAGIN 64 1 1 76160 0.475
-KAGIN 256 1 1 1091072 0.732
-KAGIN 16 1 3 10048 0.658
-KAGIN 32 1 3 32384 0.659
-KAGIN 64 1 3 113920 0.792
-KAGIN 256 1 3 1635328 2.159
-KAGIN 16 4 1 11696 0.476
-KAGIN 32 4 1 37728 0.478
-KAGIN 64 4 1 132800 0.485
-KAGIN 256 4 1 1907456 1.074
-KAGIN 16 4 3 14992 0.662
-KAGIN 32 4 3 48416 0.726
-KAGIN 64 4 3 170560 1.005
-KAGIN 256 4 3 2451712 3.092
+GIN 16 NA NA 3,522 0.368
+GIN 64 NA NA 44,802 0.372
+GIN 256 NA NA 670,722 0.433
+GIN 512 NA NA 2,652,162 0.742
+KAGIN 16 1 1 6,752 0.474
+KAGIN 32 1 1 21,696 0.475
+KAGIN 64 1 1 76,160 0.475
+KAGIN 256 1 1 1,091,072 0.732
+KAGIN 16 1 3 10,048 0.658
+KAGIN 32 1 3 32,384 0.659
+KAGIN 64 1 3 113,920 0.792
+KAGIN 256 1 3 1,635,328 2.159
+KAGIN 16 4 1 11,696 0.476
+KAGIN 32 4 1 37,728 0.478
+KAGIN 64 4 1 132,800 0.485
+KAGIN 256 4 1 1,907,456 1.074
+KAGIN 16 4 3 14,992 0.662
+KAGIN 32 4 3 48,416 0.726
+KAGIN 64 4 3 170,560 1.005
+KAGIN 256 4 3 2,451,712 3.092
 4.3 Graph Regression
@@
 To produce graph representations, we use the sum operator.
-The emerging graph representations are finally fed to an MLP (for GIN) or a KAN (for KAGIN) layer which produces the output.
+The emerging graph representations are finally fed to a 2-layer MLP (for GIN) or a 2-layer KAN (for KAGIN) which produces the output.
+We set the batch size equal to $128$ for all models.
 We train each model for $1,000$ epochs by minimizing the mean absolute error (MAE).

### v3

@@
 In recent years, Graph Neural Networks (GNNs) have become the de facto tool for learning node and graph representations.
-Most GNNs typically consist of a sequence of neighborhood aggregation (a.k.a., message passing) layers.
-Within each of these layers, the representation of each node is updated from an aggregation and transformation of its neighbours representations at the previous layer.
-The upper bound for the expressive power of message passing GNNs was reached through the use of MLPs as a transformation, due to their universal approximation capabilities.
-However, MLPs suffer from well-known limitations, which recently motivated the introduction of Kolmogorov-Arnold Networks (KANs).
-KANs rely on the Kolmogorov-Arnold representation theorem, rendering them a promising alternative to MLPs.
-In this work, we compare the performance of KANs against that of MLPs in graph learning tasks.
+Most GNNs typically consist of a sequence of neighborhood aggregation (a.k.a., message-passing) layers, within which the representation of each node is updated based on those of its neighbors.
+The most expressive message-passing GNNs can be obtained through the use of the sum aggregator and of MLPs for feature transformation, thanks to their universal approximation capabilities.
+However, the limitations of MLPs recently motivated the introduction of another family of universal approximators, called Kolmogorov-Arnold Networks (KANs) which rely on a different representation theorem.
+In this work, we compare the performance of KANs against that of MLPs on graph learning tasks.
+We evaluate two different implementations of KANs using two distinct base families of functions, namely B-splines and radial basis functions.
 We perform extensive experiments on node classification, graph classification and graph regression datasets.
-Our preliminary results indicate that while KANs are on-par with MLPs in classification tasks, they seem to have a clear advantage in the graph regression tasks.
-Code is available at https://github.com/RomanBresson/KAGNN.
+Our results indicate that KANs are on-par with or better than MLPs on all studied tasks, making them viable alternatives, at the cost of some computational complexity.
+All code will be made available upon acceptance.
 1 Introduction
-Graphs are structural representations of information which are useful for modeling many types of data.
+Graphs are structural representations of information useful for modeling different types of data.
 They arise naturally in a wide range of application domains, and their abstract nature offers increased flexibility.
@@
 In chemo-informatics, molecules are commonly modeled as graphs, with nodes corresponding to atoms and edges to chemical bonds.
-In other settings, molecules can also be nodes, with edges capturing their ability to bond with one another.
+In other settings, molecules can also correspond to nodes, with edges capturing their ability to bond with one another.
 In many cases where graph data is available, there exist problems that cannot be solved efficiently using conventional tools (e.
@@
 Recently, Graph Neural Networks (GNNs) have been established as the dominant approach for learning on graphs [2].
-Most GNNs consist of a series of message passing layers.
-Within a message passing layer, each node updates its feature vector by aggregating the feature vectors of its neighbors and combining the emerging vector with its own representation.
+Most GNNs consist of a series of message-passing layers.
+Within a message-passing layer, each node updates its feature vector by aggregating the feature vectors of its neighbors and combining the emerging vector with its own representation.
 A lot of recent work has focused on investigating the expressive power of GNNs [3].
@@
 In this setting, a model is more expressive than another model if the former can distinguish all pairs of non-isomorphic graphs that the latter can distinguish, along with other pairs that the latter cannot [4].
-Furthermore, an equivalence has also been established between the ability of GNNs to distinguish non-isomorphic graphs and their ability to approximate permutation-invariant functions on graphs [3].
+Furthermore, an equivalence was established between the ability of GNNs to distinguish non-isomorphic graphs and their ability to approximate permutation-invariant functions on graphs [5].
 This line of work gave insights into the limitations of different models [6, 7], but also led to the development of more powerful architectures [4, 8, 9].
@@
 The theorem states that any continuous function can be approximated by an MLP with at least one hidden layer, given that this layer contains enough neurons.
-Having said that, in practice the models suffer from several limitations due to non-convex loss functions, algorithms without convergence guarantees and a notorious lack of interpretability that hinders their applicability in several domains.
 Recently, Kolmogorov-Arnold Networks (KANs) [12] have emerged as promising alternatives to MLPs.
 They are based on the Kolmogorov-Arnold representation theorem [13] which states that a continuous multivariate function can be represented by a composition and sum of a fixed number of univariate functions.
-KANs substitute the learnable weights and pre-defined activation functions of MLPs, with learnable activations based on B-splines and summations.
-The initial results demonstrate that KANs have the potential to be more accurate than MLPs in low dimensions, while simultaneously being more interpretable.
-In this paper, we present a thorough empirical comparison between GNNs that use KANs to update node representations and GNNs that utilize MLPs to that end.
-Our work is orthogonal to prior work that studies the expressive power of GNNs.
-Here, we empirically compare models that are theoretically equally expressive in terms of distinguishing non-isomorphic graphs against each other, and we study the impact of the different function approximation modules (i.
-e., KANs or MLPs) on the model’s performance.
-We evaluate the different GNN models on several standard node classification, graph classification and graph regression datasets.
+KANs substitute the learnable weights and pre-defined activation functions of MLPs with learnable activations and summations.
+The initial results demonstrate that KANs have the potential to be more accurate than MLPs in low dimensions and in settings where regularity is expected.
+In this paper, we present a thorough empirical comparison of the performance between GNNs that use KANs to update node representations and GNNs that utilize MLPs to that end.
+Our intuition is that KANs could exploit the expected regularity underlying the message-passing paradigm to provide comparatively good performance on graph tasks.
+Our work is orthogonal to prior work that studies the expressive power of GNNs since we empirically compare models that are theoretically equally expressive in terms of distinguishing non-isomorphic graphs against each other, and we study the impact of the different function approximation modules (i.
+e., KANs or MLPs) on the models’ performance.
+We evaluate the different GNN models on several standard node classification, graph classification, and graph regression datasets.
 The rest of this paper is organized as follows.
-Section 2, provides an overview of the tasks we address in this paper, as well as a description of message passing GNNs and Kolmogorov-Arnold networks.
+Section 2, provides an overview of the tasks we address in this paper, as well as a description of message-passing GNNs and Kolmogorov-Arnold networks.
 In Section 3, we introduce the KAGIN (Kolmogorov-Arnold Graph Isomorphism Network) and KAGCN (Kolmogorov-Arnold Graph Convolution Network) models, which are variants of existing GNNs, and which leverage KANs to update node features within each layer.
@@
 2 Background
-2.1 Considered Graph Learning Tasks
+2.1 Graph Learning Tasks
 Before presenting the tasks on which we focus in this study, we start by introducing some key notation for graphs.
@@
 Let $g\colon V\rightarrow[n]$ denote a bijective mapping from the space of nodes to set $[n]$.
-Let $\mathcal{N}(v)$ denote the the neighbourhood of vertex $v$, i.
+Let $\mathcal{N}(v)$ denote the neighbourhood of vertex $v$, i.
 e., the set $\{u\mid\{v,u\}\in E\}$.
 The degree of a vertex $v$ is $\deg(v)=|\mathcal{N}(v)|$.
-Each node $v\in V$ is associated with a $d$-dimensional feature vector $\mathbf{x}_{v}\in\mathbb{R}^{d}$, and the feature matrix for all nodes is represented as $\mathbf{X}\in\mathbb{R}^{n\times d}$.
+Each node $v\in V$ is associated with a feature vector $\mathbf{x}_{v}\in\mathbb{R}^{d}$, and the feature matrix for all nodes is represented as $\mathbf{X}\in\mathbb{R}^{n\times d}$.
 Thus, $\mathbf{x}_{v}$ is equal to the $g(v)$-th row of $\mathbf{X}$.
 In node classification, each node $v\in V$ is associated with a label $y_{v}$ that represents a class.
-The task is to learn a function that maps nodes to their class labels, i.
-e., to learn a function $f_{\text{node}}$ such that $f_{\text{node}}(v,G,\mathbf{X})=y_{v}$.
+The task is to learn how to map nodes to their class labels, i.
+e., to learn a function $f_{\texttt{node}}$ such that $f_{\texttt{node}}(v,G,\mathbf{X})=y_{v}$.
 In graph regression/classification, the dataset consists of a collection of $N$ graphs $G_{1},\ldots,G_{N}$ along with their class labels/targets $y_{G_{1}},\ldots,y_{G_{N}}$.
 The task is then to learn a function that maps graphs to their class labels/targets, i.
-e., to learn a function $f_{\text{graph}}$ such that $f_{\text{graph}}(G,\mathbf{X})=y_{G}$, which can be discrete or continuous, for graph classification or graph regression, respectively.
+e., a function $f_{\texttt{graph}}$ such that $f_{\texttt{graph}}(G,\mathbf{X})=y_{G}$, which can be discrete or continuous, for graph classification or graph regression, respectively.
 The standard approach for learning such predictors (both for node- and graph-level tasks) is to first embed the nodes of the graph(s) into some vector space.
-That is, we aim to learn $\mathbf{H}=\texttt{EMBEDDING}(G,\textbf{X})\in\mathbb{R}^{n\times d_{e}}$ where $d_{e}$ denotes the embedding dimension.
+That is, we aim to learn $\mathbf{H}=f_{\texttt{embedding}}(G,\textbf{X})\in\mathbb{R}^{n\times d_{e}}$ where $d_{e}$ denotes the embedding dimension.
 Then, the $g(v)$-th row of matrix $\mathbf{H}$ represents the embedding of node $v$.
@@
 For node-level tasks, we can use $\mathbf{h}_{v}$ to predict directly the class label/target of node $v$.
-For graph-level tasks, we also need to apply a readout function on all the representations of the graph’s nodes to obtain a representation $\mathbf{h}_{G}=\texttt{READOUT}(\mathbf{H})$ for the entire graph.
+For graph-level tasks, we also need to apply a readout function on all the representations of the graph’s nodes to obtain a representation $\mathbf{h}_{G}=f_{\texttt{readout}}(\mathbf{H})$ for the entire graph.
 One particularly desirable property of such models is permutation invariance.
 That is, the embedding $\mathbf{h}_{G}$ of a graph needs to be the same regardless of the ordering of its nodes.
-Indeed, these orderings do not hold any semantic meaning and different orderings give rise to isomorphic graphs.
-Permutation invariance is achieved at the readout step by utilizing a permutation invariant operation over the rows of $\mathbf{H}$, such as the sum, max or mean operators.
+Indeed, these orderings do not hold any semantic meaning, and different orderings give rise to isomorphic graphs.
+Permutation invariance is achieved at the readout step by utilizing a permutation invariant operation over the rows of $\mathbf{H}$, such as the sum, max, or mean operators.
 2.2 Graph Neural Networks
-One of the most widely-used paradigms for designing such permutation invariant models is the message passing framework [12] which consists of a sequence of layers and whithin each layer the embedding of each node is computed as a learnable function of its neighbors’ embeddings.
+One of the most widely-used paradigms for designing such permutation invariant models is the message-passing framework [1] which consists of a sequence of layers and within each layer the embedding of each node is computed as a learnable function of its neighbors’ embeddings.
 Formally, the embedding $\mathbf{h}_{v}^{(\ell)}\in\mathbb{R}^{d_{\ell}}$ at layer $\ell$ is computed as follows:
-$\mathbf{h}_{v}^{(\ell)}=\phi^{(l)}\left(\mathbf{h}_{v}^{(\ell-1)},\bigoplus\limits_{u\in\mathcal{N}(v)}\mathbf{h}_{u}^{(\ell-1)}\right)$ (1)
+$\mathbf{h}_{v}^{(\ell)}=\phi^{(\ell)}\left(\mathbf{h}_{v}^{(\ell-1)},\bigoplus\limits_{u\in\mathcal{N}(v)}\mathbf{h}_{u}^{(\ell-1)}\right)$ (1)
 where $\bigoplus$ is a permutation-invariant aggregation function (e.
@@
 g., linear transformation, MLP) that combines and transforms the node’s previous embedding with the aggregated vector of its neighbors.
-As discussed above, in this paper, we focus on the functions that different GNN models employ to update node representations.
-Many existing GNNs use a $1$-layer perceptron (i.
+As discussed above, we focus here on the functions that different GNN models employ to update node representations.
+Some GNNs use a $1$-layer perceptron (i.
 e., a linear mapping followed by a non-linear activation function) within each neighborhood aggregation layer to update node features [14, 15, 16].
-For instance, each layer of the the Graph Convolutional Network (GCN) [20] is defined as follows:
+For instance, each layer of the Graph Convolutional Network (GCN) [15] is defined as follows:
 $\mathbf{h}_{v}^{(\ell)}=\sigma\left(\mathbf{W}^{(\ell)}\sum\limits_{u\in\mathcal{N}(v)\cup\{v\}}\frac{\mathbf{h}_{u}^{(\ell-1)}}{\sqrt{(\deg(v)+1)(\deg(u)+1)}}\right)$ (2)
 where $\sigma$ is a non-linear activation and $\mathbf{W}^{(\ell)}$ is a trainable weight matrix.
-However, the 1-layer perceptron is not a universal approximator of multiset functions [38].
-Thus, the emerging GNN might not be expressive enough for some tasks.
+However, the 1-layer perceptron is not a universal approximator of multiset functions [6], limiting the expressivity of the model.
 Thus, more recent models use MLPs instead of 1-layer perceptrons to update node representations [17, 18, 19, 7].
-It is well-known that standard message passing GNNs are bounded in expressiveness by the Weisfeiler-Leman (WL) test of isomorphism [38].
+It is well-known that standard message-passing GNNs are bounded in expressiveness by the Weisfeiler-Leman (WL) test of isomorphism [6].
 While two isomorphic graphs will always be mapped to the same representation by such a GNN, some non-isomorphic graphs might also be assigned identical representations.
@@
 The GIN model can achieve its full potential if proper weights (i.
-e., for the different $\texttt{MLP}^{(\ell)}$ layers and $\epsilon^{(\ell)}$) are learned.
-However, in practice, GIN might fail to learn those weights due to limited training data and due to limitations of the employed training algorithm (e.
-g., stochastic gradient descent).
-This has motivated a series of works which focused on improving the training procedure of GNNs.
+e., for the different $\texttt{MLP}^{(\ell)}$ layers and $\epsilon^{(\ell)}$) are learned, which is not guaranteed.
+This has motivated a series of works that focused on improving the training procedure of GNNs.
 For example, Ortho-GConv is an orthogonal feature transformation that can address GNNs’ unstable training [20].
-Other works have studied how to initialize the weights of the MLPs of the message passing layers of GNNs.
-It was shown that by adopting the weights of converged MLPs as the weights of corresponding GNNs can lead to performance improvements in node classification tasks [14].
-On the other hand, there exist settings where there is no need of complex learning models.
+Other works have studied how to initialize the weights of the MLPs of the message-passing layers of GNNs.
+It was shown that adopting the weights of converged MLPs as the weights of corresponding GNNs can lead to performance improvements in node classification tasks [21].
+On the other hand, there exist settings where there is no need for complex learning models.
 This has led to the development of methods for simplifying GNNs.
-This can be achieved by removing the nonlinearities between the neighborhood aggregation layers and collapsing the resulting function into a single linear transformation [35] or by feeding the node features into a neural network which generates predictions and then propagate those predictions via a personalized PageRank scheme [9].
+This can be achieved by removing the nonlinearities between the neighborhood aggregation layers and collapsing the resulting function into a single linear transformation [22] or by feeding the node features into a neural network which generates predictions and then propagating those predictions via a personalized PageRank scheme [23].
 2.3 Kolmogorov-Arnold Networks
+2.3.1 Principle
 Presented as an alternative to the MLP, the Kolmogorov-Arnold Network (KAN) architecture has recently attracted a lot of attention in the machine learning community [12].
@@
 $f(\mathbf{x})=\sum_{i=1}^{2d+1}\Phi_{i}\left(\sum_{j=1}^{d}\phi_{ij}(\mathbf{x}_{j})\right)$ (4)
-where all $\Phi_{\square}$ and $\phi_{\square}$ functions are univariate, and the sum is the only multivariate operator.
-Equation (4) can be seen as a two-step process.
+where all $\Phi_{\square}$ and $\phi_{\square}$ denote univariate functions, and the sum is the only multivariate operator.
+Equation 4 can be seen as a two-step process.
 First, a different set of univariate non-linear activation functions is applied to each dimension of the input, and then the output of those functions are summed up.
@@
 $\mathbf{x}_{i}^{\prime}=\sum\limits_{j=1}^{d}\phi_{ij}(\mathbf{x}_{j})$ (5)
-Stacking two such layers, one with input dimension $d$ and output dimension $12d+$ 1, and another with input dimension $12d+$ 1 and output dimension $1$, we obtain Equation (4), and the derived model is a universal function approximator.
+Stacking two such layers, one with input dimension $d$ and output dimension $12d+$ 1, and another with input dimension $12d+$ 1 and output dimension $1$, we obtain Equation 4, and the derived model is a universal function approximator.
 This seemingly offers a complexity advantage compared to MLPs, since the number of univariate functions required to represent any multivariate function from $\left[0,1\right]^{d}$ to $\mathbb{R}^{d^{\prime}}$ is at most $(2d^{2}+d)\times d^{\prime}$, whereas the universal approximation theorem for the MLP requires a possibly infinite number of neurons.
-However, as stated by the original paper, the behavior of such univariate functions might be arbitrarily complex (e.
-g., fractal, non-smooth), thus leading to them being non-representable, and non-learnable.
-MLPs relax the arbitrary-width constraint by stacking finite-width layers.
-Likewise, KANs relax the arbitrary-complexity constraints on the non-linearities by stacking KAN layers.
+However, as stated in the original paper, the behavior of such univariate functions might be arbitrarily complex (e.
+g., fractal, non-smooth), thus leading to them being non-representable and non-learnable.
+MLPs relax the infinite-width constraint by stacking finite-width layers.
+Likewise, KANs relax the arbitrary complexity constraints on the non-linearities by stacking KAN layers of lower complexity.
 Thus, the output of a function is given by:
-$y=\texttt{KAN}(\mathbf{x})=\Phi_{L}\circ\Phi_{L-1}\circ\cdots\circ\Phi_{1}(\mathbf{x})$ (6)
-where $\Phi_{1},\ldots,\Phi_{L}$ are KAN layers.
-The original paper uses splines (i.
+$y=\texttt{KAN}(\mathbf{x})=\Phi_{L}\circ\Phi_{L-1}\circ\cdots\circ\Phi_{1}(\mathbf{x})~{}~{}\texttt{ where $\Phi_{1},\ldots,\Phi_{L}$ are KAN layers.}$. (6)
+2.3.2 KAN Implementations and Applications
+The original paper that introduced KANs uses splines (i.
 e., trainable piecewise-polynomial functions) as nonlinearities.
@@
 The degree used for each spline (called spline order), as well as the number of splines used for each function (called grid size) are both hyperparameters of the architecture.
+Nonetheless, KANs allow for any function to serve as its base element.
+Thus, alternative implementations use different bases, such as radial-basis functions [24], which in particular alleviate some computational bottleneck of splines.
 Even though KANs were introduced very recently, they have already been applied to different problems such as in the task of satellite image classification [25] and for predicting the pressure and flow rate of flexible electrohydrodynamic pumps [26].
@@
 For instance, KANs have been evaluated in the satellite traffic forecasting task [28].
-Furthermore, they have been combined with architectures that are traditionally leveraged in time series forecasting tasks such as the Long Short-Term Memory Network [11] and the Transformer [10].
+Furthermore, they were combined with architectures usually leveraged in time series forecasting tasks such as the LSTM Network [29] and the Transformer [30].
 The work closest to ours is the one reported in [31], where the authors propose FourierKAN-GCF.
@@
 Since KAN is also a universal function approximator, we could achieve the same expressive power using KANs in lieu of MLPs.
-We thus propose the KAGIN model which is defined as follows:
+We thus propose the KAGIN model, defined as:
 $\mathbf{h}^{(\ell)}_{v}=\texttt{KAN}^{(l)}\left((1+\epsilon)\cdot\mathbf{h}^{(\ell-1)}_{v}+\sum_{u\in\mathcal{N}(v)}\mathbf{h}^{(\ell-1)}_{u}\right)$ (7)
 With theoretically-sound KANs (i.
-e., with arbitrarily complex components), this architecture is exactly as expressive as the vanilla GIN model with arbitrary layer width.
-While this is not guaranteed with the spline-based implementation with limited grid size, the empirical results in the original paper demonstrate the great expressive power of KANs [23], especially of small models and setting where regularity is expected.
+e., with arbitrarily complex components), this architecture is exactly as expressive as the vanilla GIN model with infinite layer width.
+In practice, we employ two different families of functions to serve as the base components, leading to two different layers: BS-KAGIN, based on B-splines, as in the original paper; and RBF-KAGIN, based on radial basis functions (RBFs), which were proposed as a more computationally-efficient alternative.
+While these do not guarantee universal approximation, the empirical results in the original paper demonstrate the great expressive power of KANs [12], motivating our empirical study.
 3.2 The KAGCN Layer
@@
 To this end, we also propose a variant of the GCN model.
-Specifically, we substitute the parameters and ReLU function of the standard GCN [20] model with a single KAN layer (defined in Equation (6)) to obtain the KAGCN layer:
+Specifically, we substitute the parameters and ReLU function of the standard GCN [15] model with a single KAN layer (defined in Equation 6) to obtain the KAGCN layer:
 $\mathbf{h}^{(\ell)}_{v}=\Phi^{(\ell)}\left(\sum\limits_{u\in\mathcal{N}(v)\cup\{v\}}\frac{\mathbf{h}_{u}^{(\ell-1)}}{\sqrt{(\deg(v)+1)(\deg(u)+1)}}\right)$ (8)
 where $\Phi^{(\ell)}$ denotes a single KAN layer.
-In the familiar matrix formulation, where $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ is the adjacency matrix with self-loops and $\tilde{\mathbf{D}}$ the diagonal degree matrix of $\tilde{\mathbf{A}}$, the node update rule of KAGCN can be written as:
-$\mathbf{H}^{(\ell)}=\Phi^{(\ell)}\Big{(}\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}\mathbf{H}^{(\ell-1)}\Big{)}$ (9)
-where the different rows of $\mathbf{H}^{(\ell)}$ store the representations of the different nodes of the graph.
+In the familiar matrix formulation, where $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ is the adjacency matrix with self-loops and $\tilde{\mathbf{D}}$ the diagonal degree matrix of $\tilde{\mathbf{A}}$, the node update rule of KAGCN can be written as $\mathbf{H}^{(\ell)}=\Phi^{(\ell)}\Big{(}\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}\mathbf{H}^{(\ell-1)}\Big{)}$ where the different rows of $\mathbf{H}^{(\ell)}$ store the representations of the different nodes of the graph.
+We propose both a BS-KAGCN and a RBF-KAGCN, where $\Phi$ is based respectively on B-splines and RBFs.
 4 Empirical Evaluation
-In this section, we compare the KAGIN and KAGCN models with the GIN and GCN models in the following tasks: node classification, graph classification and graph regression.
-The code for reproducing the results is available at https://github.com/RomanBresson/KAGNN.
+In this section, we compare both KAGIN and KAGCN models against the GIN and GCN models in the following tasks: node classification, graph classification, and graph regression.
+The code for reproducing the results is available at anonymized.
 All models are implemented with PyTorch [32].
-For KAN layers, we rely on a publicly available implementation [^1] https://github.com/Blealtan/efficient-kan.
+For KAN layers, we rely on publicly available implementations for B-splines  [^1] https://github.com/Blealtan/efficient-kan and for RBF  [^2] https://github.com/ZiyaoLi/fast-kan.
+For a given task, all models have consistent architectures, differing as little as possible.
 4.1 Node classification
-Datasets.
+Datasets
 To evaluate the performance of GNNs with KAN layers in the context of node classification, we use $7$ well-known datasets of varying sizes and types, including homophilic (Cora, Citeseer [15] and Ogbn-arxiv [33]) and heterophilic (Cornell, Texas, Wisconsin, Actor) networks.
-The homophilic networks are already split into training, validation and test sets, while the heterophilic datasets are accompanied by fixed $10$-fold cross validation indices.
+The homophilic networks are already split into training, validation, and test sets, while the heterophilic datasets are accompanied by fixed $10$-fold cross-validation indices.
 Experimental setup.
-For every dataset and model, we tune the values of the hyperparameters.
-Specifically, we choose the values that perform best on the validation set (i.
-e., lowest validation error).
-To find these values, we use the Optuna package [1].
-We set the number of iterations of Optuna equal to $100$ trials.
-For all models, the learning rate is chosen from $[10^{-3},10^{-2}]$, the number of message passing layers from $\{1,2,3,4\}$, the hidden dimension size from $\{8,9,\ldots,128\}$ and the weight decay from $\{0,0.0005\}$.
-For KAGIN and KAGCN, we also choose the grid size from $\{3,4,5\}$ and the spline order from $\{1,2,3,4\}$.
+We perform two types of experiments.
+First, we compare our KAGCN model on the homophilic datasets to the GCN proposed in the state of the art approach in [34].
+The results are shown in Table 1.
+One can observe that the classical GCN outperforms the KAN-based model in 2 out of 3 datasets, albeit with a narrow difference and confidence overlaps in all of the comparisons.
+Note that current setting is built to favor the classical GNNs, as it relies on techniques such as dropout, jumping knowledge, batch normalization and preliminary linear transformations.
+Hence, although the MLP-based GCN enjoys a plethora of assistive techniques examined in the past decades, the KAN-based counterpart exhibits a close performance.
+We should note here that these assistive techniques have not been developed yet for KAN-based architectures, and our results aim at motivating the development of such techniques.
+Table 1: Average classification accuracy ($\pm$ standard deviation) and model parameters based on Luo et al.[34] using Fastkan, KAGCN, GCN on the Cora, CiteSeer, and Arxiv datasets.
+Model Cora CiteSeer Arxiv
+Accuracy (%) Params Accuracy (%) Params Accuracy (%) Params
+RBF-KAGCN 81.72 $\pm$ 0.79 3,714,217 68.16 $\pm$ 0.61 6,258,808 71.39 $\pm$ 0.28 813,647
+BS-KAGCN 80.84 $\pm$ 1.61 3,702,144 70.04 $\pm$ 1.66 4,317,952 70.58 $\pm$ 0.20 1,010,432
+GCN 81.60 $\pm$ 0.40 3,262,983 71.60 $\pm$ 0.40 6,221,830 71.74 $\pm$ 0.29 2,591,784
+Table 2: Hyperparameter ranges for node classification
+Model Learning Rate Hidden Layers Hidden Dim. Grid Size Spl. Order
+GIN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[8,1024\right]$ NA NA
+BS-KAGIN/KAGCN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[2,64\right]$ $\left[1,8\right]$ $\left[1,4\right]$
+RBF-KAGIN/KAGCN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[2,64\right]$ $\left[2,32\right]$ NA
+Our second approach compares the vanilla versions of the neural architectures when we replace the MLP with KAN, as delineated in sections 3.
+We only use message passing layers, skip connections and an output layer.
+We do not utilize dropout, batchnorm, jumping knowledge etc.
+The reason we constrain ourselves to the bare minimum is to isolate and clarify the effect of KAN on the model’s performance.
+In other words, we aspire to compare the classical MLP-based GNNs as if they were in the same nascent stage as their KAN-based counterparts.
+Having said that, we perform an auxiliary experiment afterwards, to position the KAGCN in the SOTA and clarify its current utility.
+We then evaluate the $10$ models on the test set and report the average accuracy.
+For the heterophilic datasets, we report the average accuracy across the $10$ folds.
+In order for all models to exploit the same quantity of information, we fix the number of message-passing layer dataset-wise, with $2$ for Cora and CiteSeer, $3$ for Texas, Cornell, Wisconsin, and Ogbn-arxiv and $4$ for Actor.
+For every dataset and model, we tune the values of the hyperparameters using the Optuna package [35].
+For each parameterization, we train $5$ models and select the parameterization that yields the lowest mean validation loss (cross-entropy) across all $5$ models.
+We set the number of iterations of Optuna equal to $100$ trials with a TPE Sampler.
+We use early stopping with a patience of $50$.
 Once the best hyperparameter values are found, we evaluate the models on the test set.
-For the homophilic networks, we initialize and train $10$ different models (we use $10$ different random seeds).
-We then evaluate the $10$ models on the test set and report the average accuracy.
-For the heterophilic datasets, we tune each model’s hyperparameters within each fold and we report the average accuracy across the $10$ folds.
+For the homophilic networks, we initialize and train $10$ different models with the best hyperparameters (using $10$ different random seeds).
+The ranges for hyperparameters are different for each model.
+Indeed, for the same number of ”neurons”, a KAN-based layer will have a much larger number of parameters than an MLP-based layer (see Appendix 5.2 for examples).
+In order to avoid giving an unfair advantage to KAN-based layers, we thus allowed the MLP layers to be much larger, so that the largest possible model in all cases had around the same number of parameters.
+The hyperparameters ranges are given in table 2.
 Table 3: Average classification accuracy ($\pm$ standard deviation) of the KAGIN, GIN, KAGCN, and GCN models on the $7$ node classification datasets.
-Dataset KAGIN GIN KAGCN GCN
-Cora 0.7620 $\pm$ 0.0077 0.5528 $\pm$ 0.1112
-0.7826
-$\pm$ 0.0177
-0.7566 $\pm$ 0.0718
-Citeseer 0.6837 $\pm$ 0.0117 0.4748 $\pm$ 0.0574 0.6409 $\pm$ 0.0185
-0.6891
-$\pm$ 0.0074
-Ogbn-arxiv 0.2450 $\pm$ 0.0806 0.2110 $\pm$ 0.0628
-0.2997
-$\pm$ 0.0274
-0.2777$\pm$ 0.0699
-Cornell
-0.5243
-$\pm$ 0.0813
-$0$$\pm$ 0.1414 0.4892 $\pm$ 0.0547 0.4162 $\pm$ 0.0737
-Texas 0.5973 $\pm$ 0.0505
-0.6054
-$\pm$ 0.0516
-0.5811 $\pm$ 0.0440 0.5892 $\pm$ 0.0415
-Wisconsin
-0.5510
-$\pm$ 0.0720
-0.4961 $\pm$ 0.1532 0.5196 $\pm$ 0.1020 0.5216 $\pm$ 0.0729
-Actor
-0.2890
-$\pm$ 0.0139
-0.2817 $\pm$ 0.0110 0.2755 $\pm$ 0.0134 0.2852 $\pm$ 0.0129
+Model Cora Citeseer Ogbn-arxiv Cornell Texas Wisconsin Actor
+GIN 67.48$\pm$6.53 61.04$\pm$6.42 35.68$\pm$12.9 41.70$\pm$1.70 45.14$\pm$6.32 43.24$\pm$2.68 26.43$\pm$0.86
+BS-KAGIN 76.28$\pm$0.98 63.42$\pm$2.06 30.53$\pm$5.59 48.92$\pm$1.08 62.30$\pm$0.23 50.78$\pm$1.54 28.00$\pm$0.37
+RBF-KAGIN 72.34$\pm$3.74 44.26$\pm$12.0 21.77$\pm$13.2 47.49$\pm$1.85 59.92$\pm$1.88 51.73$\pm$1.51 28.91$\pm$0.40
+GCN 75.00$\pm$4.11 62.35$\pm$3.87 40.60$\pm$6.75 41.27$\pm$1.96 55.32$\pm$3.27 52.02$\pm$1.12 28.64$\pm$0.24
+BS-KAGCN 78.99$\pm$0.80 66.81$\pm$1.25 52.67$\pm$4.07 48.73$\pm$1.58 58.70$\pm$0.52 56.18$\pm$1.09 27.07$\pm$0.27
+RBF-KAGCN 63.59$\pm$7.10 48.00$\pm$8.32 54.56$\pm$3.99 45.86$\pm$2.10 58.43$\pm$1.21 53.35$\pm$1.38 27.10$\pm$0.31
 Results.
 The results are given in Table 3.
-We can see that KAGIN outperforms GIN on all but one datasets.
-On some datasets, the difference in performance between the two models is significant.
-For example, on Citeseer and Cora, KAGIN offers a respective absolute improvement of $20.89\%$ and $20.9\%$ in accuracy over GIN.
-On the other hand, KAGCN outperforms GCN on $3/7$ considered datasets, albeit being the best in $2/3$ homophilic networks and exhibiting very similar accuracy in $3/4$ heterophilic ones.
-The KAGIN and GIN models exhibit the best performance in heterophilic networks while the KAGCN and GCN models achieve the highest accuracy on the homophilic.
-The use of KANs in the GIN architecture brings performance improvements on $6/7$ datasets, but $5/7$ experiments overlap in the confidence intervals, an overlap that is more prevalent with GCN-based models in $6/7$.
-Overall we can contend that KAN has a more positive impact on the GIN architecture than on the GCN architecture.
-Moreover, the introduction of KAN does not alleviate the inherent disadvantage of the model e.g.
-GCN accuracy with heterophilic networks [41], however, it improves the models in the majority of the results i.e.
-$5/7$.
-With regards to the optimal hyperparameters, we found that KAGIN required a larger grid size on average compared to KAGCN.
-In general, a grid size of $4$ was chosen in $8/14$ experiments and the predominant spline order was $1$, appearing in $8/14$ experiments while $3$ appears $2$ times and $4$ only once.
-The number of hidden layers and their sizes varied significantly through datasets but overall KAGIN had a substantially larger average hidden layer size i.
-e., $75.4$ compared to $48$ for KAGIN, as $128$ was chosen for $3$ datasets compared to only $1$ for KAGCN.
-This is in contrast to previous findings contending that GIN-based models requires more complex learning procedures, a pattern that seems to withstand with KANs.
-Training times.
-We present in Table 2 the training time per epoch for different configurations of the KAGIN and GIN models.
-We observe that for a given number of message passing layers and hidden dimension size, the KAGIN model is computationally more expensive than GIN.
-If the grid size and spline order hyperparameters of KAGIN are set to $1$, the difference in running time between the two models is very small (less than $0.02$ seconds per epoch for all configurations).
-However, the complexity of KAGIN increases are the grid size and spline order increase.
-Overall, our results suggest that the running time of KAGIN is slightly greater than that of GIN, and by no means prohibitive.
-Table 2: Training time per epoch on the Ogbn-arxiv dataset.
-Architecture
-Message Passing
-Layers
-Hidden
-Dimension
-Grid Size Spline Order # Parameters
-Training time
-(s/epoch)
-GIN 2 16 NA NA 5,336 0.035
-GIN 2 64 NA NA 27,368 0.041
-GIN 4 16 NA NA 6,936 0.037
-GIN 4 64 NA NA 52,200 0.056
-KAGIN 2 16 1 1 31,232 0.046
-KAGIN 2 64 1 1 63,488 0.048
-KAGIN 2 16 4 1 54,656 0.059
-KAGIN 2 16 4 3 70,272 0.146
-KAGIN 4 16 1 1 38,400 0.056
-KAGIN 4 64 1 1 116,736 0.072
-KAGIN 4 64 4 1 204,288 0.109
-KAGIN 4 64 4 3 262,656 0.298
+We see that KAN-based variants beat MLP-based models on all datasets, while the B-splines-based variant outperforms both other models on $5/7$ datasets.
+We also observe that RBF-KAGIN seems to exhibit a high instability, especially on Citeseer and Ogbn-arxiv, where its performance is much lower than that of other models, and its variance a lot higher.
+We notice that the simpler GCN-based models are more efficient than the more complex GIN-based models on all three homophilic datasets, while the opposite is true on heterophilic datasets.
+This is coherent with the inherent complexity discrepency between both families of datasets [36].
+Overall, B-splines are more effective for both graph learning models, the difference being clear mostly in homophilic networks.
+Overall we contend that KAN has a positive impact on both, the GIN and the GCN architecture.
+Moreover, the introduction of KAN does not alleviate the inherent disadvantage of the model e.
+g., GCN’s low accuracy on heterophilic networks, however, it improves the models in the majority of the datasets.
 4.2 Graph Classification
 Datasets.
-In this set of experiments, we compare the KAGIN model against GIN on standard graph classification benchmark datasets [25].
-We experiment with the 7 following datasets: (1) MUTAG, (2) DD, (3) NCI1, (4) PROTEINS, (5) ENZYMES, (6) IMDB-B, (7) IMDB-M.
+In this set of experiments, we compare the KAGIN models against GIN on standard graph classification benchmark datasets [37].
+We experiment with the following $7$ datasets: (1) MUTAG, (2) DD, (3) NCI1, (4) PROTEINS, (5) ENZYMES, (6) IMDB-B, (7) IMDB-M.
 The first $5$ datasets come from bio- and chemo-informatics, while the last $2$ are social interaction datasets.
@@
 We follow the experimental protocol proposed in [38].
-Thus, we perform $10$-fold cross-validation to obtain an estimate of the generalization performance of each method, while within each fold a model is selected based on a $90\%/10\%$ split of the training set.
-We use the splits provided in [8].
-We use the Optuna package to select the model that achieves the lowest validation error.
-We set the number of iterations of Optuna equal to $100$.
-For a fair comparison, we set the number of message passing layers of both models to a fixed value for each dataset.
-Based on preliminary experiments, on MUTAG, PROTEINS, IMDB-B and IMDB-M, we set the number of layers to $2$.
+We kept our architecture simple.
+First, a number of message-passing layers extract node representation, whose model is either an MLP (resp.
+a KAN).
+We then use sum pooling to obtain graph representations.
+Finally, an MLP (resp.
+a KAN) with same architecture as those found in the message-passing layer, is used as a classifier.
+A softmax is used for predicting class probabilities.
+Thus, we perform $10$-fold cross-validation, while within each fold a model is selected based on a $90\%/10\%$ split of the training set.
+We use the pre-defined splits provided in [38].
+We use the Optuna package to select the model that achieves the lowest validation cross-entropy loss over $100$ iterations.
+For a fair comparison, we set the number of message-passing layers of all architectures to a fixed value for each dataset.
+On MUTAG, PROTEINS, IMDB-B and IMDB-M, we set the number of layers to $2$.
 On DD, we set it to $3$.
 On ENZYMES, we set it to $4$ and finally, on NCI1, we set it to $5$.
-To produce graph representations, we use the sum operator.
-The produced graph representations are fed to an MLP (for GIN) or a KAN (for KAGIN) layer which computes the output.
-We train each model for $1,000$ epochs by minimizing the cross entropy loss.
+We train each model for $1,000$ epochs (early stopping with a patience of $20$ epochs) by minimizing the cross entropy loss.
 We use the Adam optimizer for model training [39].
-We apply batch normalization [17] and dropout [33] to the output of each message passing layer.
-We also use early stopping with a patience of $20$ epochs.
-For both models, we choose the number of hidden layers from $\{2,3,4\}$ and the dropout rate from $[0.0,0.5]$.
-For GIN, we choose the hidden dimension size from $\{8,9,\ldots,256\}$ and the learning rate from $[10^{-5},10^{-2}]$.
-For KAGIN, we choose the hidden dimension size from $\{2,3,\ldots,128\}$, the grid size from $\{1,2,\ldots,16\}$, the spline order from $\{1,2,\ldots,8\}$ and the learning rate from $[10^{-4},10^{-2}]$.
-For the IMDB-B and IMDB-M datasets, where there are no node features, we annotate nodes with one-hot encodings of their degrees, up to $35$ (all degrees above $35$ are set equal to $35$).
-Once the best hyperparemeters are found for a given split, we train $3$ different models on the training set of the split (in order to limit the impact of random initialization) and evaluate them on the test set of the split.
-This yields $3$ test accuracies, and we compute their average.
-Table 3: Average classification accuracy ($\pm$ standard deviation) of the KAGIN and GIN models on the $6$ graph classification datasets.
-MUTAG DD NCI1 PROTEINS
-GIN 85.09 $\pm$ 5.97 73.68 $\pm$ 3.89 79.57 $\pm$ 1.27 70.26 $\pm$ 3.62
-KAGIN
-85.45
-$\pm$ 7.38
-75.46
-$\pm$ 4.16
-79.72
-$\pm$ 1.79
-71.78
-$\pm$ 3.12
-ENZYMES IMDB-B IMDB-M
-GIN
-53.72
-$\pm$ 9.23
-73.03 $\pm$ 2.86
-49.78
-$\pm$ 3.38
-KAGIN 42.94 $\pm$ 5.69
-73.83
-$\pm$ 4.35
-49.78
-$\pm$ 4.69
+We apply batch normalization [40] and dropout [41] to the output of each message-passing layer.
+The ranges for hyperparameters search are given in table 4.
+Once the best hyperparameters are found for a given split, we train $3$ different models on the training set of the split and evaluate them on the test set of the split.
+This yields $3$ test accuracies for this split, and we compute their average to obtain the performance for this split.
+Model Learning Rate Hidden Layers Hidden Dim. Dropout Grid Size Spl. Order
+GIN $\left[1e-4,1e-2\right]$ $\left[1,4\right]$ $\left[2,512\right]$ $\left[0,0.9\right]$ NA NA
+BS-KAGIN/KAGCN $\left[1e-4,1e-2\right]$ $\left[1,4\right]$ $\left[2,64\right]$ $\left[0,0.9\right]$ $\left[2,32\right]$ $\left[1,4\right]$
+RBF-KAGIN/KAGCN $\left[1e-4,1e-2\right]$ $\left[1,4\right]$ $\left[2,64\right]$ $\left[0,0.9\right]$ $\left[2,16\right]$ NA
+Table 4: Hyperparameter ranges for node classification
+Table 5: Average classification accuracy ($\pm$ standard deviation) of the KAGIN and GIN models on the $7$ graph classification datasets.
+MUTAG DD NCI1 PROTEINS ENZYMES IMDB-B IMDB-M
+GIN 85.26 $\pm$ 7.30 74.78 $\pm$ 4.65 78.32 $\pm$ 2.49 71.01 $\pm$ 2.81 53.28 $\pm$ 1.23 73.13 $\pm$ 3.97 50.27 $\pm$ 4.01
+BS-KAGIN 85.08 $\pm$ 6.07 75.37 $\pm$ 3.37 79.44 $\pm$ 1.67 74.09 $\pm$ 3.60 42.72 $\pm$ 8.94 73.37 $\pm$ 4.03 49.91 $\pm$ 3.47
+RBF-KAGIN 85.48 $\pm$ 5.36 75.66 $\pm$ 3.08 78.33 $\pm$ 1.81 73.29 $\pm$ 3.56 41.17 $\pm$ 8.32 71.97 $\pm$ 3.81 48.76 $\pm$ 4.53
+GCN 56.34 $\pm$ 15.1 63.24 $\pm$ 3.82 60.42 $\pm$ 1.01 60.46 $\pm$ 2.06 22.06 $\pm$ 5.58 68.70 $\pm$ 3.63 46.24 $\pm$ 4.83
+BS-KAGCN 75.55 $\pm$ 8.14 73.31 $\pm$ 4.45 71.59 $\pm$ 3.31 75.55 $\pm$ 3.62 57.67 $\pm$ 8.03 70.83 $\pm$ 3.95 50.00 $\pm$ 4.41
+RBF-KAGCN 73.52 $\pm$ 11.6 68.30 $\pm$ 4.53 63.95 $\pm$ 7.80 74.18 $\pm$ 3.64 59.39 $\pm$ 8.03 72.60 $\pm$ 4.61 48.53 $\pm$ 4.03
 Results.
-Table 3 illustrates the average classification accuracies and the corresponding standard deviations of the two models on the different datasets.
-We observe that the two models achieve similar levels of performance.
-KAGIN outperforms GIN on $5$ out of the $7$ datasets.
-However, the difference in performance between the two models is very small.
-This suggests that the two models are similar in terms of expressive power.
-On ENZYMES, however, KAGIN was found to perform much worse than GIN.
-Note that ENZYMES consists of more classes ($6$ classes in total) than the rest of the datasets, while ENZYMES and PROTEINS are the only datasets where the nodes of the graphs are annotated with continuous features (on the rest of the datasets, they are annotated with one-hot encodings).
-We hypothesize that the difference in performance is due to the inability of KANs to handle those continuous features.
-We thus normalized the node features within each fold by removing the mean of each feature (computed from the training samples) and then dividing by the corresponding standard deviation.
-We re-conducted the experiment and the average accuracy increased by approximately $6\%$ ($48.77\%$ instead of $42.94\%$).
-Therefore, it turns out that in some settings it might be harder for KAN layers to handle continuous features than for MLPs.
-Training times.
-We give in Table 4 an overview of the training times for different GIN and KAGIN architectures.
-We provide the number of parameters of each architecture.
-We notice that, for the same number of parameters, KAN is slower than its MLP counterpart.
-This is particularly sensitive to grid size and spline order.
-This makes sense since, using splines, each parameter involves more complex computations than the usual multiplication/summation of traditional MLP neurons.
-Moreover, some of this performance difference might come from how optimized the implementation is.
-It would be an interesting next step to study the relation of size to performance for KANs and MLPs, since, intuitively, the expressivity of splines should allow for smaller networks.
-Table 4: Training time per epoch on NCI1 dataset (i.
-e., only the forward and backward pass, looped over all batches, averaged over $50$ epochs).
-All compared models consist of $5$ message passing layers, and of $2$ hidden layers in their respective MLP/KAN.
-Batch size is set to $128$.
-Architecture
-Hidden
-Dimension
-Grid Size Spline Order # Parameters
-Training time
-(s/epoch)
-GIN 16 NA NA 3,522 0.368
-GIN 64 NA NA 44,802 0.372
-GIN 256 NA NA 670,722 0.433
-GIN 512 NA NA 2,652,162 0.742
-KAGIN 16 1 1 6,752 0.474
-KAGIN 32 1 1 21,696 0.475
-KAGIN 64 1 1 76,160 0.475
-KAGIN 256 1 1 1,091,072 0.732
-KAGIN 16 1 3 10,048 0.658
-KAGIN 32 1 3 32,384 0.659
-KAGIN 64 1 3 113,920 0.792
-KAGIN 256 1 3 1,635,328 2.159
-KAGIN 16 4 1 11,696 0.476
-KAGIN 32 4 1 37,728 0.478
-KAGIN 64 4 1 132,800 0.485
-KAGIN 256 4 1 1,907,456 1.074
-KAGIN 16 4 3 14,992 0.662
-KAGIN 32 4 3 48,416 0.726
-KAGIN 64 4 3 170,560 1.005
-KAGIN 256 4 3 2,451,712 3.092
+Table 5 illustrates the average classification accuracies and the corresponding standard deviations of the three models on the different datasets.
+We observe that GIN and both KAGIN  models achieve similar levels of performance on all datasets except ENZYMES.
+In particular, RBF-KAGIN  and BS-KAGIN  both have the best accuracy on $2$ datasets each, and GIN on $1$ dataset.
+Regarding the GCN variants, we observe that the KAN-based variants now have a significant advantage over the linear GCN, confirming that a single KAN layer is more expressive than a simple fully-connexted layer.
+BS-KAGCN  and RBF-KAGCN  obtain the best overall accuracy on $1$ dataset each, while the vanilla GCN performs often significantly poorly.
+These experiments show that, on a graph-level task, which requires significantly more power than node-level tasks, KAN-based models seem to perform better than their MLP-based counterparts.
+While this difference is not significant among GIN-like architectures, it appears that KAN-based GCNs benefit from the higher expressivity in a single layer, and attain performances much closer to their GIN-like counterparts, sometimes even outperforming them.
 4.3 Graph Regression
@@
 ZINC-12K consists of $12,000$ molecules.
-The task is to predict the constrained solubility of molecules, an important chemical property for designing generative GNNs for molecules.
+The task is to predict the constrained solubility of molecules, an important property for designing generative GNNs for molecules.
 The dataset is already split into training, validation and test sets ($10,000$, $1,000$ and $1,000$ graphs in the training, validation and test sets, respectively).
 QM9 contains approximately $134,000$ organic molecules.
-Each molecule consists of Hydrogen (H), Carbon (C), Oxygen (O), Nitrogen (N), and Flourine (F) atoms and contain up to 9 heavy (non Hydrogen) atoms.
+Each molecule consists of Hydrogen, Carbon, Oxygen, Nitrogen, and Flourine atoms and contain up to 9 heavy (non Hydrogen) atoms.
 The task is to predict $12$ target properties for each molecule.
-The dataset was divided into a training, a validation and a test set according to a $80\%/10\%/10\%$ split.
+The dataset was divided into a training, a validation, and a test set according to a $80\%/10\%/10\%$ split.
 Experimental setup.
 We perform grid search to select values for the different hyperparameters.
-For both models, we choose the number of hidden layers from $\{2,3,4\}$, and the learning rate from $\{10^{-3},10^{-4}\}$.
+We choose the number of hidden layers from $\{2,3,4\}$ for GIN and $\{1,2,3\}$ for KAN.
+We use different ranges for each model was because we wanted to limit the size advantage of KAN against MLP, since a layer with the same width of KAN has more parameters than an MLP one).
+For both models we search a learning rate from $\{10^{-3},10^{-4}\}$.
 For GIN, we choose the hidden dimension size from $\{32,64,128,256,512,1024\}$, while for KAGIN, we choose it from $\{4,8,16,32,64,128,256\}$.
-For KAGIN, we also select the grid size from $\{1,3,5,8,10\}$ and the spline order from $\{3,5\}$.
+For KAGIN, we also select the grid size from $\{1,3,5,8,10\}$ and for BS-KAGIN the spline order from $\{3,5\}$.
 To produce graph representations, we use the sum operator.
@@
 We also use early stopping with a patience of $20$ epochs.
-For ZINC-12K, we also use an embedding layer which maps node features into $100$-dimensional vectors.
+For ZINC-12K, we also use an embedding layer that maps node features into $100$-dimensional vectors.
 We choose the configuration that achieves the lowest validation error.
 Once the best configuration is found, we run $10$ experiments and report the average performance on the test set.
-For both datasets and models, we set the number of message passing layers to $4$.
+For both datasets and models, we set the number of message-passing layers to $4$.
 On QM9, we performed a joint regression of the $12$ targets.
-Table 5: Average MAE ($\pm$ standard deviation) of the KAGIN and GIN models on the ZINC-12K and QM9 datasets.
+Table 6: Average MAE ($\pm$ standard deviation) of the KAGIN and GIN models on graph regression.
 ZINC-12K QM9
 GIN 0.4131 $\pm$ 0.0215 0.0969 $\pm$ 0.0017
-KAGIN
+BS-KAGIN
 0.3000
@@
 $\pm$ 0.0007
+RBF-KAGIN 0.3026 $\pm$ 0.0663 0.0778 $\pm$ 0.0023
 Results.
 The results are shown in Table 6.
-We observe that on both considered datasets, KAGIN significantly outperforms the GIN model.
-Note that these datasets are significantly larger (in terms of number of samples) compared to the graph classification datasets of Table 3.
-More specifically, KAGIN offers an absolute improvement of approximately $0.11$ and $0.03$ in MAE over GIN.
+We observe that on both considered datasets, both KAGIN variants significantly outperform the GIN model.
+Note that these datasets are significantly larger (in terms of the number of samples) compared to the graph classification datasets of Table 5.
+The results also indicate that BS-KAGIN outperforms RBF-KAGIN on both ZINC-12K and QM9.
+However, the difference in performance between the two models is very small.
+Thus, it appears that both B-splines and RBFs are good candidates as KANs’ base components in regression tasks.
+In addition, we should note that the BS-KAGIN model offers an absolute improvement of approximately $0.11$ and $0.03$ in MAE over GIN.
 Those improvements suggest that KANs might be more effective than MLPs in regression tasks.
+5 Analysis of Size and Time
+5.1 Selected models
+Following the above results, and it is legitimate to wonder whether the performance of the KAN models comes from a larger size than their MLP-based counterparts.
+We show in this section that it is not the case.
+Using the graph-classification models trained for the results provided in 4.2, we display here three relations: the size-to-performance relation, the size-to-training time relation and the training time-to-performance relation.
+Figures 2 and 4 show these relations for datasets NCI1 and IMDB-MULTI respectively.
+We also show in 8 and 8 the numerical values associated to these graphs.
+The rest of the graph classification datasets are provided in App.
+5.2.
+A first thing that we can see from both figures is that, for the same parameter count, B-splines-based KANs are significantly slower than both other architectures.
+RBF-based KANs are much faster, showing that this implementation succeeds in alleviating the computational bottleneck induced by the sequential computation necessary for B-splines.
+Finally, MLPs are the fastest for equal parameter counts.
+While these results were expected, we notice in both the tables and figures that RBF-KAGINmodels end up significantly smaller than their MLP-based counterparts, while maintaining equal or better performance, and about equal training time.
+On the other hand, the B-Splines based method is significantly slower, by sometimes a 5 fold factor.
+This establishes the RBF-based architecture as the best compromise between scalability and performance, although further work could be made in studying the performance of very small KAN-based against MLP-based models.
+Figure 1: Parameters count/size/performance comparison, GIN, NCI1
+Figure 2: Parameters count/size/performance comparison, GCN, NCI1
+Figure 3: Parameters count/size/performance comparison, GIN, IMDB-MULTI
+Figure 4: Parameters count/size/performance comparison, GCN, IMDB-MULTI
+Table 7: Time and parameter sizes for selected models, NCI1
+Model Accuracy Parameter count time per epoch (s)
+BS-KAGIN 0.794 $\pm$ 0.016 572338.4 2.220
+RBF-KAGIN 0.783 $\pm$ 0.017 139168.0 0.358
+GIN 0.783 $\pm$ 0.024 1337369.8 0.359
+BS-KAGCN 0.716 $\pm$ 0.031 90196.5 0.752
+RBF-KAGCN 0.640 $\pm$ 0.075 166960.5 0.356
+GCN 0.604 $\pm$ 0.046 379634.0 0.277
+Table 8: Time and parameter sizes for selected models, IMDB-MULTI
+Model Accuracy Parameter count time per epoch (s)
+BS-KAGIN 0.499 $\pm$ 0.033 296418.8 0.367
+RBF-KAGIN 0.488 $\pm$ 0.043 48705.1 0.169
+GCN 0.503 $\pm$ 0.038 330987.6 0.146
+BS-KAGIN 0.500 $\pm$ 0.042 46052.3 0.216
+RBF-KAGIN 0.485 $\pm$ 0.038 40118.5 0.161
+MLP 0.462 $\pm$ 0.046 111911.4 0.143
+5.2 Theoretical models
+We give in Table 9 the training time per epoch for different configurations of the BS-KAGIN, RBF-KAGIN and GIN models on a graph-level task.
+We observe that for a given number of message-passing layers and hidden dimension size, GIN is the fastest, followed by RBF-KAGIN.
+BS-KAGIN is noticeably slower than the others.
+It suffers doubly: its parameter count increases much quicker than both other models for same hyperparameters; but even for similar numbers of parameters, BS-KAGIN is by far the slowest.
+Spline order seems to have a larger impact on running time than grid size.
+This shows that RBFs do alleviate some of the computational bottlenecks of the B-splines from the original KAN paper.
+Table 16 in Appendix C shows similar observations on a node-level task.
+Table 9: Training time per epoch on NCI1 dataset. $5$ message-passing layers. batch size set to $128$.
+Architecture
+Hidden
+Dimension
+Hidden
+Layers
+Grid Size Spline Order # Parameters
+Training time
+(s/epoch)
+GIN 256 2 NA NA 670722 0.303
+GIN 256 4 NA NA 1465346 0.390
+GIN 512 2 NA NA 2652162 0.541
+GIN 512 4 NA NA 5814274 0.951
+RBF-KAGIN 256 1 1 NA 822872 0.386
+RBF-KAGIN 256 1 4 NA 1639274 0.599
+RBF-KAGIN 256 3 1 NA 3191408 0.783
+RBF-KAGIN 256 3 4 NA 6367142 1.577
+RBF-KAGIN 512 1 1 NA 3218520 0.807
+RBF-KAGIN 512 1 4 NA 6424170 1.441
+RBF-KAGIN 512 3 1 NA 12674160 2.198
+RBF-KAGIN 512 3 4 NA 25317030 4.429
+BS-KAGIN 256 1 1 1 1091072 0.446
+BS-KAGIN 256 1 1 4 1907456 2.964
+BS-KAGIN 256 1 4 1 1907456 0.720
+BS-KAGIN 256 1 4 4 2723840 4.240
+BS-KAGIN 256 3 1 1 4236800 0.937
+BS-KAGIN 256 3 1 4 7412480 9.703
+BS-KAGIN 256 3 4 1 7412480 1.963
+BS-KAGIN 256 3 4 4 10588160 14.572
+6 Discussion
+Prominent application domains.
+KANs, both in our work and in the recent literature, seem to be a promising alternative to MLPs for regression tasks.
+They rely on combinations of continuous functions which could explain their improved performance for continuous outputs.
+Moreover, this property and their empirical performance in regression tasks renders them an ideal candidate for physics-inspired architectures [44], which address problems of high importance, from climate forecasting [45] to atomic simulations [46] and rely heavily on graph-based architectures.
+Explainability matters.
+The potential of explainability for KAN-based models is arguably better compared to that of MLPs with post-hoc methods (e.g.
+Shapley values, gradient-based), since closed-form symbolic formulas can be drawn from the former, providing complete overview of the underlying computations.
+Hence, KAN-based graph models could potentially prove more friendly to practitioners from application fields, such as biology and physics, who tend to be skeptical of the black-box approach.
+GNNs for fixed graphs inspired from KANs.
+Each KAN architecture consists of a set of neurons.
+Each one of those neurons is connected to other neurons.
+Thus, KANs can be naturally modeled as graphs where nodes represent neurons and edges represent the connections between them.
+Each layer of a KAN is actually a complete bipartite graph, while a KAN has learnable activation functions on the edges of its associated graph.
+There actually exist datasets that contain multiple graphs which share the same graph structure, but differ in the node and edge features.
+Two examples of such datasets are the MNIST and CIFAR10 graph classification datasets [47].
+Since the graph structure is fixed (and identical for all graphs), we could associate each edge of the graph with one or more learnable activation functions (shared across graphs), similar to what KANs do.
+This would allow the model to learn different functions for each edge of the graph, leading potentially to high-quality node and graph representations.
+Advantages and future work.
+Finally, we discuss potential advantages that KANs might have over MLPs, and leave their investigation for future work.
+First, their ability to accurately fit smooth functions could prove relevant on datasets where variables interact with some regular patterns.
+Second, their interpretability could be leveraged to provide explanations on learned models, giving insights into the nature of interactions among entities.
+Finally, a thorough study of the effect of the different hyperparameters could be pursued, allowing to fully exploit the splines or radial basis functions while retaining small networks.
 7 Conclusion
@@
 We then compared those architectures against each other in both node- and graph-level tasks.
-In the classification tasks, there does not appear to be a clear winner, with both architectures outperforming each other on some datasets.
-In the graph regression task, however, preliminary results seem to indicate that KAN has an advantage over MLP.
-This paper shows, through its preliminary results, that such KAN-based GNNs are valid alternatives to the traditional MLP-based models.
-We thus believe that these models deserve the attention of the graph machine-learning community.
-Finally, we discuss potential advantages that KANs might have over MLPs, and leave their investigation for future work.
-First, their ability to accurately fit smooth functions could prove highly relevant on datasets where variables interact with some regular patterns.
-Second, their interpretability could be leveraged to provide explanations on learned models, giving insights into the nature of interactions among entities.
-Finally, a thorough study of the effect of the different hyperparameters could be leveraged, allowing to fully exploit the richness of splines while retaining small networks.
-6 Acknowledgements
-This work was partially supported by the The Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation and partner Swedish universities and industry.
+For graph and node classification, there does not appear to be a clear winner.
+For graph regression, our results indicate that KAN has an advantage over MLP, with the B-spline variant proving superior in terms of performance.
+Nonetheless, this gain comes at the cost of computational efficiency.
+In this regard, the RBF-based vairant offers a slightly lesser performance improvement, but proves much more efficient in terms of training time.
+This paper shows that such KAN-based GNNs are valid alternatives to the traditional MLP-based models.
+We thus believe that these models deserve the attention of the graph machine-learning community, and that future work should be carried out to identify their specific advantages and weaknesses, allowing for an enlightened choice between KANs and MLPs based on the task and context at hand.
+Appendix A Further Graph Classification Experiments
+Besides the graph classification experiments that are presented in the main paper, we also experimented with the Peptides-func dataset, a graph classification dataset from the Long Range Graph Benchmark [48].
+The dataset consists of $15,535$ graphs where the average number of nodes per graph is equal to $150.94$.
+We evaluated GIN, BS-KAGIN and RBF-KAGIN on this dataset.
+We performed a grid search where for all models we chose the learning rate from $\{0.001,0.0001\}$, the number of hidden layers of MLPs and KANs from $\{1,2,3\}$ and the hidden dimension from $\{64,128\}$.
+We set the number of neighborhood aggregation layers to $5$ for all models.
+For both BS-KAGIN and RBF-KAGIN, grid size was chosen from $\{3,5\}$, and for BS-KAGIN, spline order was chosen from $\{3,5\}$.
+Each experiment was run $4$ times with $4$ different seeds and Table 10 illustrates the average precision achieved by the three models.
+Table 10: Average Precision ($\pm$ standard deviation) of the KAGIN and GIN models on the Peptides-func dataset.
+Peptides-func
+GIN 55.01 $\pm$ 0.75
+BS-KAGIN
+58.71
+$\pm$ 1.02
+RBF-KAGIN 57.68 $\pm$ 2.05
+We observe that BS-KAGIN and RBF-KAGIN outperform GIN on the Peptides-func dataset.
+Both models offer an absolute improvement of more than $2.5\%$ in average precision over GIN.
+Thus, even in tasks where long-range interactions between nodes need to be captured, for a fixed receptive field, KAN layers can help the models better capture those interactions.
+Appendix B Time/Performance/Accuracy analysis
+Figures 6 to 14 and tables 11 to 15 regarding the time/size/accuracy analysis on the graph classification task.
+Table 11: Time and parameter sizes for selected models, IMDB-BINARY
+Model Accuracy Parameter count time per epoch (s)
+BS-KAGIN 0.734 $\pm$ 0.038 148236.3 0.225
+RBF-KAGIN 0.720 $\pm$ 0.036 168383.6 0.129
+GIN 0.731 $\pm$ 0.038 1172520.5 0.110
+BS-KAGCN 0.708 $\pm$ 0.037 39053.5 0.150
+RBF-KAGCN 0.726 $\pm$ 0.044 37322.5 0.109
+GCN 0.687 $\pm$ 0.034 118826.3 0.099
+Table 12: Time and parameter sizes for selected models, ENZYMES
+Model Accuracy Parameter count time per epoch (s)
+BS-KAGIN 0.427 $\pm$ 0.085 450136.6 0.336
+RBF-KAGIN 0.412 $\pm$ 0.079 115346.8 0.061
+GIN 0.533 $\pm$ 0.117 1623033.2 0.048
+BS-KAGCN 0.577 $\pm$ 0.076 83636.3 0.103
+RBF-KAGCN 0.594 $\pm$ 0.053 125854.4 0.050
+GCN 0.221 $\pm$ 0.053 280040.0 0.039
+Table 13: Time and parameter sizes for selected models, DD
+Model Accuracy Parameter count time per epoch (s)
+BS-KAGIN 0.754 $\pm$ 0.032 169417.4 1.206
+RBF-KAGIN 0.757 $\pm$ 0.029 147005.1 0.236
+GIN 0.748 $\pm$ 0.044 168957.4 0.144
+BS-KAGCN 0.733 $\pm$ 0.042 110278.6 0.596
+RBF-KAGCN 0.683 $\pm$ 0.043 110899.2 0.185
+GCN 0.632 $\pm$ 0.036 342023.2 0.225
+Table 14: Time and parameter sizes for selected models, MUTAG
+Model Accuracy Parameter count time per epoch (s)
+BS-KAGIN 0.851 $\pm$ 0.058 236235.8 0.056
+RBF-KAGIN 0.855 $\pm$ 0.051 62530.0 0.020
+GIN 0.853 $\pm$ 0.069 848213.0 0.014
+BS-KAGCN 0.755 $\pm$ 0.077 19496.1 0.022
+RBF-KAGCN 0.735 $\pm$ 0.110 47769.4 0.017
+GCN 0.563 $\pm$ 0.143 72072.2 0.012
+Table 15: Time and parameter sizes for selected models, PROTEINS_full
+Model Accuracy Parameter count time per epoch (s)
+BS-KAGIN 0.741 $\pm$ 0.034 321367.7 0.472
+RBF-KAGIN 0.733 $\pm$ 0.034 51498.8 0.096
+MLP 0.710 $\pm$ 0.027 942781.0 0.073
+BS-KAGCN 0.756 $\pm$ 0.034 37688.7 0.133
+RBF-KAGCN 0.742 $\pm$ 0.035 48950.4 0.078
+GCN 0.605 $\pm$ 0.020 8199.8 0.061
+Figure 5: Parameters count/size/performance comparison, GIN, IMDB-BINARY
+Figure 6: Parameters count/size/performance comparison, GCN, IMDB-BINARY
+Figure 7: Parameters count/size/performance comparison, GIN, ENZYMES
+Figure 8: Parameters count/size/performance comparison, GCN, ENZYMES
+Figure 9: Parameters count/size/performance comparison, GIN, DD
+Figure 10: Parameters count/size/performance comparison, GCN, DD
+Figure 11: Parameters count/size/performance comparison, GIN, MUTAG
+Figure 12: Parameters count/size/performance comparison, GCN, MUTAG
+Figure 13: Parameters count/size/performance comparison, GIN, PROTEINS_full
+Figure 14: Parameters count/size/performance comparison, GCN, PROTEINS_full
+Appendix C Training Times - Node Classification
+Table 16: Training time per epoch on the node-classification task (Cora dataset, 3 MP-layers).
+Architecture
+Hidden
+Dimension
+Hidden
+Layers
+Grid Size Spline Order # Parameters
+Training time
+(s/epoch)
+GCN 256 NA NA NA 378934 0.003
+GCN 1024 NA NA NA 1485622 0.005
+BS-KAGCN 256 NA 1 1 1514684 0.017
+BS-KAGCN 256 NA 1 4 2650697 0.160
+BS-KAGCN 256 NA 4 1 2650697 0.031
+BS-KAGCN 256 NA 4 4 3786710 0.222
+BS-KAGCN 512 NA 1 1 2989244 0.020
+BS-KAGCN 512 NA 1 4 5231177 0.179
+BS-KAGCN 512 NA 4 1 5231177 0.039
+BS-KAGCN 512 NA 4 4 7473110 0.254
+RBF-KAGCN 256 NA 1 NA 1142524 0.016
+RBF-KAGCN 256 NA 4 NA 2278543 0.031
+RBF-KAGCN 1024 NA 1 NA 4462588 0.029
+RBF-KAGCN 1024 NA 4 NA 8916367 0.057
+GIN 256 2 NA NA 867335 0.007
+GIN 256 4 NA NA 1130503 0.008
+GIN 1024 2 NA NA 5042183 0.020
+GIN 1024 4 NA NA 9240583 0.033
+BS-KAGIN 256 1 1 1 1514684 0.017
+BS-KAGIN 256 1 1 4 2650697 0.158
+BS-KAGIN 256 1 4 1 2650697 0.032
+BS-KAGIN 256 1 4 4 3786710 0.224
+BS-KAGIN 256 3 1 1 3990528 0.025
+BS-KAGIN 256 3 1 4 6983424 0.229
+BS-KAGIN 256 3 4 1 6983424 0.051
+BS-KAGIN 256 3 4 4 9976320 0.334
+BS-KAGIN 512 1 1 1 2989244 0.021
+BS-KAGIN 512 1 1 4 5231177 0.184
+BS-KAGIN 512 1 4 1 5231177 0.041
+BS-KAGIN 512 1 4 4 7473110 0.261
+BS-KAGIN 512 3 1 1 10078208 0.046
+BS-KAGIN 512 3 1 4 17636864 0.354
+BS-KAGIN 512 3 4 1 17636864 0.089
+BS-KAGIN 512 3 4 4 25195520 0.499
+RBF-KAGIN 256 1 1 NA 1142524 0.017
+RBF-KAGIN 256 1 4 NA 2278543 0.032
+RBF-KAGIN 256 3 1 NA 3002487 0.023
+RBF-KAGIN 256 3 4 NA 5995401 0.046
+RBF-KAGIN 1024 1 1 NA 4462588 0.03
+RBF-KAGIN 1024 1 4 NA 8916367 0.057
+RBF-KAGIN 1024 3 1 NA 21429879 0.091
+RBF-KAGIN 1024 3 4 NA 42838665 0.185

### v4

@@
 In this work, we compare the performance of KANs against that of MLPs on graph learning tasks.
+We implement three new KAN-based GNN layers, inspired respectively by the GCN, GAT and GIN layers.
 We evaluate two different implementations of KANs using two distinct base families of functions, namely B-splines and radial basis functions.
-We perform extensive experiments on node classification, graph classification and graph regression datasets.
-Our results indicate that KANs are on-par with or better than MLPs on all studied tasks, making them viable alternatives, at the cost of some computational complexity.
-All code will be made available upon acceptance.
+We perform extensive experiments on node classification, link prediction, graph classification and graph regression datasets.
+Our results indicate that KANs are on-par with or better than MLPs on all tasks studied in this paper.
+We also show that the size and training speed of RBF-based KANs is only marginally higher than for MLPs, making them viable alternatives.
+Code available at https://github.com/RomanBresson/KAGNN.
 1 Introduction
@@
 g., graph algorithms) and require the use of machine learning techniques.
-For instance, in the field of chemo-informatics, the standard approach for estimating the quantum mechanical properties of molecules leverages computationally expensive density functional theory computations [1].
+For instance, in the field of chemo-informatics, the standard approach for estimating the quantum mechanical properties of molecules leverages computationally expensive density functional theory computations (Gilmer et al., 2017).
 Machine learning methods could serve as a more efficient alternative to those methods.
-Recently, Graph Neural Networks (GNNs) have been established as the dominant approach for learning on graphs [2].
+Recently, Graph Neural Networks (GNNs) have been established as the dominant approach for learning on graphs (Wu et al., 2020).
 Most GNNs consist of a series of message-passing layers.
 Within a message-passing layer, each node updates its feature vector by aggregating the feature vectors of its neighbors and combining the emerging vector with its own representation.
-A lot of recent work has focused on investigating the expressive power of GNNs [3].
+A lot of recent work has focused on investigating the expressive power of GNNs (Li and Leskovec, 2022).
 There exist different definitions of expressive power, however, the most common definition is concerned with the number of pairs of non-isomorphic graphs that a GNN model can distinguish.
 Two graphs are isomorphic if there exists an edge-preserving bijection between their respective sets of nodes.
-In this setting, a model is more expressive than another model if the former can distinguish all pairs of non-isomorphic graphs that the latter can distinguish, along with other pairs that the latter cannot [4].
-Furthermore, an equivalence was established between the ability of GNNs to distinguish non-isomorphic graphs and their ability to approximate permutation-invariant functions on graphs [5].
-This line of work gave insights into the limitations of different models [6, 7], but also led to the development of more powerful architectures [4, 8, 9].
-Most maximally-expressive GNN models rely on multi-layer perceptrons (MLPs) as their main building blocks, due to their universal approximation capabilities [10, 11].
+In this setting, a model is more expressive than another model if the former can distinguish all pairs of non-isomorphic graphs that the latter can distinguish, along with other pairs that the latter cannot (Morris et al., 2019).
+Furthermore, an equivalence was established between the ability of GNNs to distinguish non-isomorphic graphs and their ability to approximate permutation-invariant functions on graphs (Chen et al., 2019).
+This line of work gave insights into the limitations of different models (Xu et al., 2019; Nikolentzos et al., 2020), but also led to the development of more powerful architectures (Morris et al., 2019; Maron et al., 2019; Morris et al., 2020a).
+Most maximally-expressive GNN models rely on multi-layer perceptrons (MLPs) as their main building blocks, due to their universal approximation capabilities (Cybenko, 1989; Hornik et al., 1989).
 The theorem states that any continuous function can be approximated by an MLP with at least one hidden layer, given that this layer contains enough neurons.
-Recently, Kolmogorov-Arnold Networks (KANs) [12] have emerged as promising alternatives to MLPs.
-They are based on the Kolmogorov-Arnold representation theorem [13] which states that a continuous multivariate function can be represented by a composition and sum of a fixed number of univariate functions.
+Recently, Kolmogorov-Arnold Networks (KANs) (Liu et al., 2025) have emerged as promising alternatives to MLPs.
+They are based on the Kolmogorov-Arnold representation theorem (Kolmogorov, 1957) which states that a continuous multivariate function can be represented by a composition and sum of a fixed number of univariate functions.
 KANs substitute the learnable weights and pre-defined activation functions of MLPs with learnable activations and summations.
@@
 e., KANs or MLPs) on the models’ performance.
-We evaluate the different GNN models on several standard node classification, graph classification, and graph regression datasets.
+We implement two layers, one based on the Graph Convolutional Network (GCN) (Kipf and Welling, 2017) and one based on the Graph Isomorphism Network (GIN) (Xu et al., 2019).
+We evaluate the different GNN models on several standard node classification, link prediction, graph classification, and graph regression datasets.
 The rest of this paper is organized as follows.
 Section 3, provides an overview of the tasks we address in this paper, as well as a description of message-passing GNNs and Kolmogorov-Arnold networks.
-In Section 3, we introduce the KAGIN (Kolmogorov-Arnold Graph Isomorphism Network) and KAGCN (Kolmogorov-Arnold Graph Convolution Network) models, which are variants of existing GNNs, and which leverage KANs to update node features within each layer.
-In Section 4, we present extensive empirical results comparing the above models with their vanilla counterparts in several tasks.
-Finally, Section 7 concludes the paper.
+In Section 4, we introduce the KAGIN (Kolmogorov-Arnold Graph Isomorphism Network), KAGCN (Kolmogorov-Arnold Graph Convolution Network) and KAGAT (Kolmogorov-Arnold Graph ATtention Network) models, which are variants of existing GNNs, and which leverage KANs to update node features within each layer.
+In Section 5, we present extensive empirical results comparing the above models with their vanilla counterparts in several tasks, and in Section 6 we evaluate the time efficiency of our architectures.
+Finally, Section 7 presents perspectives for future work and Section 8 concludes the paper.
+2 Related Work
+Kolmogorov-Arnold Networks have recently gained significant attention in the machine learning community (Somvanshi et al., 2024).
+Even though KANs were introduced very recently, they have already been applied to different problems such as in the task of satellite image classification (Cheon, 2024), for predicting the pressure and flow rate of flexible electrohydrodynamic pumps (Peng et al., 2024), and for rumor detection (Jiang et al., 2024).
+So far, most efforts have focused on time series data (Xu et al., 2024a; Barašin et al., 2024).
+For instance, KANs have been evaluated in the satellite traffic forecasting task (Vaca-Rubio et al., 2024).
+Furthermore, they were combined with architectures usually leveraged in time series forecasting tasks such as LSTM networks (Genet and Inzirillo, 2024a) and the Transformer (Genet and Inzirillo, 2024b).
+Recently, KAN-based convolutional neural networks have also been proposed which consist of kernels made of learnable non-linear functions that use splines (Bodner et al., 2024).
+These neural networks have been utilized for intrusion detection (Wang et al., 2025).
+KANs have also paved their way into celebrated architectures such as PointNet (Kashefi, 2024).
+Recent work has focused on improving different aspects of KANs such as making them more parameter-efficient (Ta et al., 2025), robust to overfitting (Altarabichi, 2024), and extending them to the complex domain (Wolff et al., 2025).
+More importantly, a framework was recently introduced which allows a seamless interplay between KANs and science (Liu et al., 2024).
+Specifically, multiplication was incorporated as a built-in modularity of KANs.
+Furthermore, the work strengthened the connection between KANs and scientific problems by highlighting the identification of key features, modular structures, and symbolic formulas.
+Note that our study is not the only work in the intersection of KANs and graph learning algorithms (Kiamari et al., 2024; De Carlo et al., 2024; Zhang and Zhang, 2024; Li et al., 2024).
+However, while this paper provides an extensive evaluation across different settings and datasets, other works focus on a single task (e.
+g., node classification) or are limited to small-sized datasets.
+(Kiamari et al., 2024) propose two variants of the GCN architecture, where the feed-forward layer which updates node representations is replaced by KAN layers.
+The two variants differ from each other in how the new node representations are produced.
+The one transforms node representations before aggregation (and is thus similar to our KAGCN model), while the other after aggregation.
+However, the two variants are only evaluated on a single node classification dataset.
+(De Carlo et al., 2024) propose a neighborhood aggregation layer where spline-based activation functions are applied to the representations of the neighbors before aggregation, while another spline-based activation function is applied to the output of the aggregation.
+A final spline-based transformation then produces the final node representations.
+While the proposed model is evaluated in different tasks (i.
+e., node classification, graph classification and link prediction), all the considered datasets are small-sized.
+In another recent study, (Zhang and Zhang, 2024) also propose a variant of the GCN architecture, where a KAN is employed to transform node representations after aggregation.
+The proposed model is only evaluated on a single dataset where a graph is constructed from one-dimensional acoustic signals for axial flow pump fault diagnosis.
+Recent work has also investigated the ability of GNNs that employ KAN layers to solve problems in chemo- and bio-informatics such as the prediction of properties of molecules (Li et al., 2024) and the prediction of the binding affinity of small molecules to protein targets (Ahmed and Sifat, 2024), respectively.
+Such models were also applied to the problem of graph collaborative filtering (Xu et al., 2024b).
 3 Background
@@
 The task is to learn how to map nodes to their class labels, i.
-e., to learn a function $f_{\texttt{node}}$ such that $f_{\texttt{node}}(v,G,\mathbf{X})=y_{v}$.
+e., to learn a function $f_{\text{node}}$ such that $f_{\text{node}}(v,G,\mathbf{X})=y_{v}$.
 In graph regression/classification, the dataset consists of a collection of $N$ graphs $G_{1},\ldots,G_{N}$ along with their class labels/targets $y_{G_{1}},\ldots,y_{G_{N}}$.
 The task is then to learn a function that maps graphs to their class labels/targets, i.
-e., a function $f_{\texttt{graph}}$ such that $f_{\texttt{graph}}(G,\mathbf{X})=y_{G}$, which can be discrete or continuous, for graph classification or graph regression, respectively.
+e., a function $f_{\text{graph}}$ such that $f_{\text{graph}}(G,\mathbf{X})=y_{G}$, which can be discrete or continuous, for graph classification or graph regression, respectively.
 The standard approach for learning such predictors (both for node- and graph-level tasks) is to first embed the nodes of the graph(s) into some vector space.
-That is, we aim to learn $\mathbf{H}=f_{\texttt{embedding}}(G,\textbf{X})\in\mathbb{R}^{n\times d_{e}}$ where $d_{e}$ denotes the embedding dimension.
+That is, we aim to learn $\mathbf{H}=f_{\text{embedding}}(G,\textbf{X})\in\mathbb{R}^{n\times d_{e}}$ where $d_{e}$ denotes the embedding dimension.
 Then, the $g(v)$-th row of matrix $\mathbf{H}$ represents the embedding of node $v$.
@@
 For node-level tasks, we can use $\mathbf{h}_{v}$ to predict directly the class label/target of node $v$.
-For graph-level tasks, we also need to apply a readout function on all the representations of the graph’s nodes to obtain a representation $\mathbf{h}_{G}=f_{\texttt{readout}}(\mathbf{H})$ for the entire graph.
+For graph-level tasks, we also need to apply a readout function on all the representations of the graph’s nodes to obtain a representation $\mathbf{h}_{G}=f_{\text{readout}}(\mathbf{H})$ for the entire graph.
 One particularly desirable property of such models is permutation invariance.
@@
 3.2 Graph Neural Networks
-One of the most widely-used paradigms for designing such permutation invariant models is the message-passing framework [1] which consists of a sequence of layers and within each layer the embedding of each node is computed as a learnable function of its neighbors’ embeddings.
+One of the most widely-used paradigms for designing such permutation invariant models is the message-passing framework (Gilmer et al., 2017) which consists of a sequence of layers and within each layer the embedding of each node is computed as a learnable function of its neighbors’ embeddings.
 Formally, the embedding $\mathbf{h}_{v}^{(\ell)}\in\mathbb{R}^{d_{\ell}}$ at layer $\ell$ is computed as follows:
@@
 Some GNNs use a $1$-layer perceptron (i.
-e., a linear mapping followed by a non-linear activation function) within each neighborhood aggregation layer to update node features [14, 15, 16].
-For instance, each layer of the Graph Convolutional Network (GCN) [15] is defined as follows:
+e., a linear mapping followed by a non-linear activation function) within each neighborhood aggregation layer to update node features (Duvenaud et al., 2015; Kipf and Welling, 2017; Zhang et al., 2018).
+For instance, each layer of the Graph Convolutional Network (GCN) (Kipf and Welling, 2017) is defined as follows:
 $\mathbf{h}_{v}^{(\ell)}=\sigma\left(\mathbf{W}^{(\ell)}\sum\limits_{u\in\mathcal{N}(v)\cup\{v\}}\frac{\mathbf{h}_{u}^{(\ell-1)}}{\sqrt{(\deg(v)+1)(\deg(u)+1)}}\right)$ (2)
 where $\sigma$ is a non-linear activation and $\mathbf{W}^{(\ell)}$ is a trainable weight matrix.
-However, the 1-layer perceptron is not a universal approximator of multiset functions [6], limiting the expressivity of the model.
-Thus, more recent models use MLPs instead of 1-layer perceptrons to update node representations [17, 18, 19, 7].
-It is well-known that standard message-passing GNNs are bounded in expressiveness by the Weisfeiler-Leman (WL) test of isomorphism [6].
+However, the 1-layer perceptron is not a universal approximator of multiset functions (Xu et al., 2019), limiting the expressivity of the model.
+Thus, more recent models use MLPs instead of 1-layer perceptrons to update node representations (Balcilar et al., 2021; Dasoulas et al., 2021; Murphy et al., 2019; Nikolentzos et al., 2020).
+It is well-known that standard message-passing GNNs are bounded in expressiveness by the Weisfeiler-Leman (WL) test of isomorphism (Xu et al., 2019).
 While two isomorphic graphs will always be mapped to the same representation by such a GNN, some non-isomorphic graphs might also be assigned identical representations.
-A model that can achieve the same expressive power as the WL test, given sufficient width and depth of the MLP, is the Graph Isomorphism Network (GIN) [6], which is defined as follows:
+A model that can achieve the same expressive power as the WL test, given sufficient width and depth of the MLP, is the Graph Isomorphism Network (GIN) (Xu et al., 2019), which is defined as follows:
 $\mathbf{h}^{(\ell)}_{v}=\texttt{MLP}^{(\ell)}\left((1+\epsilon^{(\ell)})\cdot\mathbf{h}^{(\ell-1)}_{v}+\sum_{u\in\mathcal{N}(v)}\mathbf{h}^{(\ell-1)}_{u}\right)$ (3)
@@
 This has motivated a series of works that focused on improving the training procedure of GNNs.
-For example, Ortho-GConv is an orthogonal feature transformation that can address GNNs’ unstable training [20].
+For example, Ortho-GConv is an orthogonal feature transformation that can address GNNs’ unstable training (Guo et al., 2022).
 Other works have studied how to initialize the weights of the MLPs of the message-passing layers of GNNs.
-It was shown that adopting the weights of converged MLPs as the weights of corresponding GNNs can lead to performance improvements in node classification tasks [21].
+It was shown that adopting the weights of converged MLPs as the weights of corresponding GNNs can lead to performance improvements in node classification tasks (Han et al., 2023).
 On the other hand, there exist settings where there is no need for complex learning models.
 This has led to the development of methods for simplifying GNNs.
-This can be achieved by removing the nonlinearities between the neighborhood aggregation layers and collapsing the resulting function into a single linear transformation [22] or by feeding the node features into a neural network which generates predictions and then propagating those predictions via a personalized PageRank scheme [23].
+This can be achieved by removing the nonlinearities between the neighborhood aggregation layers and collapsing the resulting function into a single linear transformation (Wu et al., 2019) or by feeding the node features into a neural network which generates predictions and then propagating those predictions via a personalized PageRank scheme (Gasteiger et al., 2019).
+Finally, the Graph Attention Network (GAT)(Veličković et al., 2017) leverages the multi-head attention mechanism among nodes to give each neighbor an adapted importance in the aggregation scheme.
+$\mathbf{h}^{(\ell)}_{v}=\sum_{u\in\mathcal{N}(v)\cup\{v\}}\alpha^{(\ell)}_{v,u}\mathbf{W}^{(\ell)}\mathbf{h}_{u}^{(\ell-1)}$ (4)
+where $\mathbf{W}^{(\ell)}$ denotes a single linear layer and
+$\alpha_{uv}^{(\ell)}=\frac{\exp\left(\texttt{LeakyReLU}\left(\left[\mathbf{W}^{(\ell)}h^{(\ell-1)}_{u}+\mathbf{W}^{(\ell)}h^{(\ell-1)}_{v}\right]\right)\right)}{\sum_{u\in\mathcal{N}(v)\cup\{v\}}\left(\exp\left(\texttt{LeakyReLU}\left(\left[\mathbf{W}^{(\ell)}h^{(\ell-1)}_{u}+\mathbf{W}^{(\ell)}h^{(\ell-1)}_{v}\right]\right)\right)\right)}$ (5)
+are the neighbor-wise attention weights for a given head.
 3.3 Kolmogorov-Arnold Networks
 3.3.1 Principle
-Presented as an alternative to the MLP, the Kolmogorov-Arnold Network (KAN) architecture has recently attracted a lot of attention in the machine learning community [12].
+Presented as an alternative to the MLP, the Kolmogorov-Arnold Network (KAN) architecture has recently attracted a lot of attention in the machine learning community (Liu et al., 2025).
 As mentioned above, this model relies on the Kolmogorov-Arnold representation theorem, which states that any multivariate function $f:\left[0,1\right]^{d}\rightarrow\mathbb{R}$ can be written as:
@@
 Thus, the output of a function is given by:
-$y=\texttt{KAN}(\mathbf{x})=\Phi_{L}\circ\Phi_{L-1}\circ\cdots\circ\Phi_{1}(\mathbf{x})~{}~{}\texttt{ where $\Phi_{1},\ldots,\Phi_{L}$ are KAN layers.}$. (6)
+$y=\texttt{KAN}(\mathbf{x})=\Phi_{L}\circ\Phi_{L-1}\circ\cdots\circ\Phi_{1}(\mathbf{x})~{}~{}\text{ where $\Phi_{1},\ldots,\Phi_{L}$ are KAN layers.}$. (8)
 3.3.2 KAN Implementations and Applications
@@
 Nonetheless, KANs allow for any function to serve as its base element.
-Thus, alternative implementations use different bases, such as radial-basis functions [24], which in particular alleviate some computational bottleneck of splines.
-Even though KANs were introduced very recently, they have already been applied to different problems such as in the task of satellite image classification [25] and for predicting the pressure and flow rate of flexible electrohydrodynamic pumps [26].
-So far, most efforts have focused on time series data [27].
-For instance, KANs have been evaluated in the satellite traffic forecasting task [28].
-Furthermore, they were combined with architectures usually leveraged in time series forecasting tasks such as the LSTM Network [29] and the Transformer [30].
-The work closest to ours is the one reported in [31], where the authors propose FourierKAN-GCF.
-This is a GNN model designed for the task of graph collaborative filtering where the feature transformation in the neighborhood aggregation layers is performed by KANs.
+Thus, alternative implementations use different bases, such as radial-basis functions (Li, 2024), which in particular alleviate some computational bottleneck of splines.
 4 KAN-based GNN Layers
-We next derive variants of the GIN and GCN models which use KANs to transform the node features instead of fully-connected layers or MLPs.
+We next derive three KAGNN layers, i.e. variants of the GIN, GCN and GAT models which use KANs to transform the node features instead of fully-connected layers or MLPs.
 4.1 The KAGIN Layer
@@
 In practice, we employ two different families of functions to serve as the base components, leading to two different layers: BS-KAGIN, based on B-splines, as in the original paper; and RBF-KAGIN, based on radial basis functions (RBFs), which were proposed as a more computationally-efficient alternative.
-While these do not guarantee universal approximation, the empirical results in the original paper demonstrate the great expressive power of KANs [12], motivating our empirical study.
+While these do not guarantee universal approximation, the empirical results in the original paper demonstrate the great expressive power of KANs (Liu et al., 2025), motivating our empirical study.
 4.2 The KAGCN Layer
@@
 To this end, we also propose a variant of the GCN model.
-Specifically, we substitute the parameters and ReLU function of the standard GCN [15] model with a single KAN layer (defined in Equation 6) to obtain the KAGCN layer:
+Specifically, we substitute the linear transformation of the standard GCN (Kipf and Welling, 2017) model with a single KAN layer (defined in Equation 8) to obtain the KAGCN layer:
 $\mathbf{h}^{(\ell)}_{v}=\Phi^{(\ell)}\left(\sum\limits_{u\in\mathcal{N}(v)\cup\{v\}}\frac{\mathbf{h}_{u}^{(\ell-1)}}{\sqrt{(\deg(v)+1)(\deg(u)+1)}}\right)$ (10)
@@
 We propose both a BS-KAGCN and a RBF-KAGCN, where $\Phi$ is based respectively on B-splines and RBFs.
+4.3 The KAGAT Layer
+Finally, we substitute the linear transformation of the standard graph attention network (GAT)  (Veličković et al., 2017) model with a single KAN layer (defined in Equation 8) to obtain the KAGAT layer:
+$\mathbf{h}^{(\ell)}_{v}=\sum_{u\in\mathcal{N}(v)\cup\{v\}}\alpha^{(\ell)}_{v,u}\Phi^{(\ell)}\left(\mathbf{h}_{u}^{(\ell-1)}\right)$ (11)
+where $\Phi^{(\ell)}$ denotes a single KAN layer and
+$\alpha_{uv}^{(\ell)}=\frac{\exp\left(\texttt{LeakyReLU}\left(\left[\Phi^{(\ell)}\left(h^{(\ell-1)}_{u}\right)+\Phi^{(\ell)}\left(h^{(\ell-1)}_{v}\right)\right]\right)\right)}{\sum_{u\in\mathcal{N}(v)\cup\{v\}}\left(\exp\left(\texttt{LeakyReLU}\left(\left[\Phi^{(\ell)}\left(h^{(\ell-1)}_{u}\right)+\Phi^{(\ell)}\left(h^{(\ell-1)}_{v}\right)\right]\right)\right)\right)}$ (12)
 5 Empirical Evaluation
-In this section, we compare both KAGIN and KAGCN models against the GIN and GCN models in the following tasks: node classification, graph classification, and graph regression.
-The code for reproducing the results is available at anonymized.
-All models are implemented with PyTorch [32].
+In this section, we compare all three KAGNN models against MLP-based GNNs on the following tasks: node classification, link prediction, graph classification, and graph regression.
+All models are implemented with PyTorch (Paszke et al., 2019).
 For KAN layers, we rely on publicly available implementations for B-splines  [^1] https://github.com/Blealtan/efficient-kan and for RBF  [^2] https://github.com/ZiyaoLi/fast-kan.
 For a given task, all models have consistent architectures, differing as little as possible.
+All GAT-based models use 4 heads.
 5.1 Node classification
 Datasets
-To evaluate the performance of GNNs with KAN layers in the context of node classification, we use $7$ well-known datasets of varying sizes and types, including homophilic (Cora, Citeseer [15] and Ogbn-arxiv [33]) and heterophilic (Cornell, Texas, Wisconsin, Actor) networks.
+To evaluate the performance of GNNs with KAN layers in the context of node classification, we use $7$ well-known datasets of varying sizes and types, including homophilic (Cora, Citeseer (Kipf and Welling, 2017) and Ogbn-arxiv (Hu et al., 2020)) and heterophilic (Cornell, Texas, Wisconsin, Actor) networks (Zhu et al., 2021).
 The homophilic networks are already split into training, validation, and test sets, while the heterophilic datasets are accompanied by fixed $10$-fold cross-validation indices.
+Thus, we clone the split for homophilic datasets $10$ times in order to have a similar number of experiments across all datasets.
 Experimental setup.
 We perform two types of experiments.
-First, we compare our KAGCN model on the homophilic datasets to the GCN proposed in the state of the art approach in [34].
+First, we compare our KAGCN model on the homophilic datasets to the GCN proposed in the state of the art approach in (Luo et al., 2024).
 The results are shown in Table 1.
 One can observe that the classical GCN outperforms the KAN-based model in 2 out of 3 datasets, albeit with a narrow difference and confidence overlaps in all of the comparisons.
-Note that current setting is built to favor the classical GNNs, as it relies on techniques such as dropout, jumping knowledge, batch normalization and preliminary linear transformations.
+Note that the current setting is built to favor the classical GNNs, as it relies on techniques such as dropout, jumping knowledge, batch normalization and preliminary linear transformations.
 Hence, although the MLP-based GCN enjoys a plethora of assistive techniques examined in the past decades, the KAN-based counterpart exhibits a close performance.
 We should note here that these assistive techniques have not been developed yet for KAN-based architectures, and our results aim at motivating the development of such techniques.
-Table 1: Average classification accuracy ($\pm$ standard deviation) and model parameters based on Luo et al.[34] using Fastkan, KAGCN, GCN on the Cora, CiteSeer, and Arxiv datasets.
+Table 1: Average classification accuracy ($\pm$ standard deviation) and model parameters based on Luo et al.(Luo et al., 2024) using GCN-based models on the Cora, CiteSeer, and Arxiv datasets.
 Model Cora CiteSeer Arxiv
 Accuracy (%) Params Accuracy (%) Params Accuracy (%) Params
+GCN 81.60 $\pm$ 0.40 3,262,983 71.60 $\pm$ 0.40 6,221,830 71.74 $\pm$ 0.29 2,591,784
+BS-KAGCN 80.84 $\pm$ 1.61 3,702,144 70.04 $\pm$ 1.66 4,317,952 70.58 $\pm$ 0.20 1,010,432
 RBF-KAGCN 81.72 $\pm$ 0.79 3,714,217 68.16 $\pm$ 0.61 6,258,808 71.39 $\pm$ 0.28 813,647
-BS-KAGCN 80.84 $\pm$ 1.61 3,702,144 70.04 $\pm$ 1.66 4,317,952 70.58 $\pm$ 0.20 1,010,432
-GCN 81.60 $\pm$ 0.40 3,262,983 71.60 $\pm$ 0.40 6,221,830 71.74 $\pm$ 0.29 2,591,784
-Table 2: Hyperparameter ranges for node classification
-Model Learning Rate Hidden Layers Hidden Dim. Grid Size Spl. Order
-GIN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[8,1024\right]$ NA NA
-BS-KAGIN/KAGCN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[2,64\right]$ $\left[1,8\right]$ $\left[1,4\right]$
-RBF-KAGIN/KAGCN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[2,64\right]$ $\left[2,32\right]$ NA
 Our second approach compares the vanilla versions of the neural architectures when we replace the MLP with KAN, as delineated in sections 4.
-We only use message passing layers, skip connections and an output layer.
-We do not utilize dropout, batchnorm, jumping knowledge etc.
-The reason we constrain ourselves to the bare minimum is to isolate and clarify the effect of KAN on the model’s performance.
-In other words, we aspire to compare the classical MLP-based GNNs as if they were in the same nascent stage as their KAN-based counterparts.
-Having said that, we perform an auxiliary experiment afterwards, to position the KAGCN in the SOTA and clarify its current utility.
-We then evaluate the $10$ models on the test set and report the average accuracy.
-For the heterophilic datasets, we report the average accuracy across the $10$ folds.
+In order to clarify the effect of KAN on the model’s performance, we only use dropout and batchnorm as additional techniques.
+Indeed, we aspire to compare the classical MLP-based GNNs as if they were in the same nascent stage as their KAN-based counterparts.
 In order for all models to exploit the same quantity of information, we fix the number of message-passing layer dataset-wise, with $2$ for Cora and CiteSeer, $3$ for Texas, Cornell, Wisconsin, and Ogbn-arxiv and $4$ for Actor.
-For every dataset and model, we tune the values of the hyperparameters using the Optuna package [35].
-For each parameterization, we train $5$ models and select the parameterization that yields the lowest mean validation loss (cross-entropy) across all $5$ models.
-We set the number of iterations of Optuna equal to $100$ trials with a TPE Sampler.
-We use early stopping with a patience of $50$.
-Once the best hyperparameter values are found, we evaluate the models on the test set.
-For the homophilic networks, we initialize and train $10$ different models with the best hyperparameters (using $10$ different random seeds).
-The ranges for hyperparameters are different for each model.
-Indeed, for the same number of ”neurons”, a KAN-based layer will have a much larger number of parameters than an MLP-based layer (see Appendix 5.2 for examples).
-In order to avoid giving an unfair advantage to KAN-based layers, we thus allowed the MLP layers to be much larger, so that the largest possible model in all cases had around the same number of parameters.
-The hyperparameters ranges are given in table 2.
-Table 3: Average classification accuracy ($\pm$ standard deviation) of the KAGIN, GIN, KAGCN, and GCN models on the $7$ node classification datasets.
+All attention-based models used 4 heads.
+For every dataset and model, we tune the values of the hyperparameters using the Optuna package (Akiba et al., 2019) with $100$ trials (parameterizations), a TPE Sampler and set early stopping patience to $50$.
+For each parameterization, and for each split, we train $3$ models in order to mitigate randomness.
+We then select the parameterization with lowest validation error across all splits and models.
+We use a different hyperparameter range for MLP-based models, B-Splines-based models and RBF-based models (see Table 8).
+Indeed, for the same number of ”neurons”, a KAN-based layer will have a much larger number of parameters than an MLP-based layer (see Appendix C.1 for examples).
+In order to avoid giving an unfair advantage to KAN-based layers, we thus allowed the MLP layers to be larger.
+Table 2: Average node classification accuracies ($\pm$ standard deviation).
 Model Cora Citeseer Ogbn-arxiv Cornell Texas Wisconsin Actor
-GIN 67.48$\pm$6.53 61.04$\pm$6.42 35.68$\pm$12.9 41.70$\pm$1.70 45.14$\pm$6.32 43.24$\pm$2.68 26.43$\pm$0.86
-BS-KAGIN 76.28$\pm$0.98 63.42$\pm$2.06 30.53$\pm$5.59 48.92$\pm$1.08 62.30$\pm$0.23 50.78$\pm$1.54 28.00$\pm$0.37
-RBF-KAGIN 72.34$\pm$3.74 44.26$\pm$12.0 21.77$\pm$13.2 47.49$\pm$1.85 59.92$\pm$1.88 51.73$\pm$1.51 28.91$\pm$0.40
-GCN 75.00$\pm$4.11 62.35$\pm$3.87 40.60$\pm$6.75 41.27$\pm$1.96 55.32$\pm$3.27 52.02$\pm$1.12 28.64$\pm$0.24
-BS-KAGCN 78.99$\pm$0.80 66.81$\pm$1.25 52.67$\pm$4.07 48.73$\pm$1.58 58.70$\pm$0.52 56.18$\pm$1.09 27.07$\pm$0.27
-RBF-KAGCN 63.59$\pm$7.10 48.00$\pm$8.32 54.56$\pm$3.99 45.86$\pm$2.10 58.43$\pm$1.21 53.35$\pm$1.38 27.10$\pm$0.31
+GCN 74.01 $\pm$ 0.81 57.73 $\pm$ 1.21 53.21 $\pm$ 0.13 66.13 $\pm$ 6.09 72.34 $\pm$ 5.92 78.63 $\pm$ 5.43 32.81 $\pm$ 1.11
+BS-KAGCN 70.46 $\pm$ 2.00 56.45 $\pm$ 1.62 53.76 $\pm$ 0.31 67.75 $\pm$ 6.54 72.16 $\pm$ 5.68 78.63 $\pm$ 5.26 34.45 $\pm$ 1.12
+RBF-KAGCN 48.06 $\pm$ 1.40 47.55 $\pm$ 1.43 53.65 $\pm$ 0.20 67.75 $\pm$ 7.24 71.62 $\pm$ 4.69 77.45 $\pm$ 4.11 34.77 $\pm$ 1.20
+GAT 76.51 $\pm$ 0.79 60.91 $\pm$ 1.04 53.68 $\pm$ 0.12 66.85 $\pm$ 5.59 69.01 $\pm$ 6.14 78.17 $\pm$ 5.70 32.51 $\pm$ 1.17
+BS-KAGAT 78.28 $\pm$ 0.83 56.87 $\pm$ 1.67 55.17 $\pm$ 0.12 67.66 $\pm$ 6.94 73.78 $\pm$ 6.39 78.43 $\pm$ 5.47 31.66 $\pm$ 1.23
+RBF-KAGAT 72.62 $\pm$ 1.64 48.65 $\pm$ 2.39 54.24 $\pm$ 0.25 64.14 $\pm$ 6.14 71.71 $\pm$ 6.49 76.34 $\pm$ 4.34 34.04 $\pm$ 0.94
+GIN 70.85 $\pm$ 1.75 47.98 $\pm$ 0.91 53.58 $\pm$ 0.74 69.10 $\pm$ 4.85 73.51 $\pm$ 5.87 78.24 $\pm$ 4.76 33.46 $\pm$ 1.38
+BS-KAGIN 65.78 $\pm$ 1.68 55.77 $\pm$ 1.51 53.05 $\pm$ 0.55 69.64 $\pm$ 7.40 72.97 $\pm$ 5.63 78.63 $\pm$ 5.16 34.26 $\pm$ 1.06
+RBF-KAGIN 72.77 $\pm$ 1.35 46.87 $\pm$ 1.93 53.56 $\pm$ 0.37 67.30 $\pm$ 6.44 73.60 $\pm$ 5.05 76.73 $\pm$ 4.88 34.37 $\pm$ 0.95
 Results.
 The results are given in Table 2.
-We see that KAN-based variants beat MLP-based models on all datasets, while the B-splines-based variant outperforms both other models on $5/7$ datasets.
-We also observe that RBF-KAGIN seems to exhibit a high instability, especially on Citeseer and Ogbn-arxiv, where its performance is much lower than that of other models, and its variance a lot higher.
-We notice that the simpler GCN-based models are more efficient than the more complex GIN-based models on all three homophilic datasets, while the opposite is true on heterophilic datasets.
-This is coherent with the inherent complexity discrepency between both families of datasets [36].
-Overall, B-splines are more effective for both graph learning models, the difference being clear mostly in homophilic networks.
-Overall we contend that KAN has a positive impact on both, the GIN and the GCN architecture.
-Moreover, the introduction of KAN does not alleviate the inherent disadvantage of the model e.
-g., GCN’s low accuracy on heterophilic networks, however, it improves the models in the majority of the datasets.
+We see that a B-splines-based variant outperforms other models on $5/7$ datasets, and both other architecture win on one dataset each.
+Nonetheless, these are often by small margins.
+We also notice poor performance of RBF-KAGCN on Cora, which we attribute to overfitting.
+Overall,there seems to be a positive impact of using KAN-based architectures.
+5.2 Link Prediction
+Datasets and Experimental Setup.
+In this setting, we focus on the task of link prediction.
+We use two datasets, Cora and CiteSeer, following the implementation and protocol provided in (Li et al., 2023; Mao et al., 2024).
+We performed a grid-search over learning rate, hidden dimension, number of hidden layers, grid size and spline order, each time training models on 10 splits.
+We then selected the model with lowest validation loss to obtain the best configuration, and reported its test ROC-AUC (over all 10 splits).
+Table 3: AUC for link prediction of the KAGIN and GIN models on the link-prediction datasets.
+Cora CiteSeer Cora CiteSeer Cora CiteSeer
+GCN 95.36 $\pm$ 0.23 96.72 $\pm$ 0.44 GAT 94.02 $\pm$ 0.48 96.10 $\pm$ 0.49 GIN 94.97 $\pm$ 0.40 95.37 $\pm$ 0.23
+BS-KAGCN 94.37 $\pm$ 0.46 94.91 $\pm$ 0.62 BS-KAGAT 94.71 $\pm$ 0.65 95.40 $\pm$ 0.33 BS-KAGIN 94.64 $\pm$ 0.44 95.75 $\pm$ 0.39
+RBF-KAGCN 94.87 $\pm$ 0.45 95.44 $\pm$ 0.67 RBF-KAGAT 94.96 $\pm$ 0.42 90.09 $\pm$ 14.1 RBF-KAGIN 95.07 $\pm$ 0.27 97.07 $\pm$ 0.40
+Results.
+Table 3 displays the results on the link prediction task.
+We see that GCN is the best performer on Cora, and RBF-KAGIN is the best on CiteSeer, albeit with often close scores among models.
+We also notice some instability of RBF-KAGAT on CiteSeer.
 5.3 Graph Classification
 Datasets.
-In this set of experiments, we compare the KAGIN models against GIN on standard graph classification benchmark datasets [37].
-We experiment with the following $7$ datasets: (1) MUTAG, (2) DD, (3) NCI1, (4) PROTEINS, (5) ENZYMES, (6) IMDB-B, (7) IMDB-M.
-The first $5$ datasets come from bio- and chemo-informatics, while the last $2$ are social interaction datasets.
+In this set of experiments, we evaluate the KAGNN models on standard graph classification benchmark datasets (Morris et al., 2020b).
+We experiment with the following $7$ datasets: MUTAG, DD, NCI1, PROTEINS, ENZYMES, IMDB-B, IMDB-M.
 Experimental setup.
-We follow the experimental protocol proposed in [38].
-We kept our architecture simple.
-First, a number of message-passing layers extract node representation, whose model is either an MLP (resp.
+We follow the experimental protocol proposed in (Errica et al., 2020).
+We kept our architecture simple: first, a number of message-passing layers extract node representation, whose model is either an MLP (resp.
 a KAN).
-We then use sum pooling to obtain graph representations.
+We then use pooling (sum-pooling for GINs and GATs, mean-pooling for GCNs) to obtain graph representations.
 Finally, an MLP (resp.
-a KAN) with same architecture as those found in the message-passing layer, is used as a classifier.
+a KAN) with same architecture as those in the message-passing layer, is used as a classifier.
 A softmax is used for predicting class probabilities.
 Thus, we perform $10$-fold cross-validation, while within each fold a model is selected based on a $90\%/10\%$ split of the training set.
-We use the pre-defined splits provided in [38].
+We use the pre-defined splits provided in (Errica et al., 2020).
 We use the Optuna package to select the model that achieves the lowest validation cross-entropy loss over $100$ iterations.
-For a fair comparison, we set the number of message-passing layers of all architectures to a fixed value for each dataset.
+For fair comparison, we fix the number of message-passing layers of all architectures for each dataset.
 On MUTAG, PROTEINS, IMDB-B and IMDB-M, we set the number of layers to $2$.
@@
 We train each model for $1,000$ epochs (early stopping with a patience of $20$ epochs) by minimizing the cross entropy loss.
-We use the Adam optimizer for model training [39].
-We apply batch normalization [40] and dropout [41] to the output of each message-passing layer.
+We use the Adam optimizer for model training (Kingma and Ba, 2015).
+We apply batch normalization (Ioffe and Szegedy, 2015) and dropout (Srivastava et al., 2014) to the output of each message-passing layer.
 The ranges for hyperparameters search are given in table 11.
@@
 This yields $3$ test accuracies for this split, and we compute their average to obtain the performance for this split.
-Model Learning Rate Hidden Layers Hidden Dim. Dropout Grid Size Spl. Order
-GIN $\left[1e-4,1e-2\right]$ $\left[1,4\right]$ $\left[2,512\right]$ $\left[0,0.9\right]$ NA NA
-BS-KAGIN/KAGCN $\left[1e-4,1e-2\right]$ $\left[1,4\right]$ $\left[2,64\right]$ $\left[0,0.9\right]$ $\left[2,32\right]$ $\left[1,4\right]$
-RBF-KAGIN/KAGCN $\left[1e-4,1e-2\right]$ $\left[1,4\right]$ $\left[2,64\right]$ $\left[0,0.9\right]$ $\left[2,16\right]$ NA
-Table 4: Hyperparameter ranges for node classification
-Table 5: Average classification accuracy ($\pm$ standard deviation) of the KAGIN and GIN models on the $7$ graph classification datasets.
+Table 4: Average graph classification accuracy ($\pm$ standard deviation).
 MUTAG DD NCI1 PROTEINS ENZYMES IMDB-B IMDB-M
-GIN 85.26 $\pm$ 7.30 74.78 $\pm$ 4.65 78.32 $\pm$ 2.49 71.01 $\pm$ 2.81 53.28 $\pm$ 1.23 73.13 $\pm$ 3.97 50.27 $\pm$ 4.01
-BS-KAGIN 85.08 $\pm$ 6.07 75.37 $\pm$ 3.37 79.44 $\pm$ 1.67 74.09 $\pm$ 3.60 42.72 $\pm$ 8.94 73.37 $\pm$ 4.03 49.91 $\pm$ 3.47
-RBF-KAGIN 85.48 $\pm$ 5.36 75.66 $\pm$ 3.08 78.33 $\pm$ 1.81 73.29 $\pm$ 3.56 41.17 $\pm$ 8.32 71.97 $\pm$ 3.81 48.76 $\pm$ 4.53
-GCN 56.34 $\pm$ 15.1 63.24 $\pm$ 3.82 60.42 $\pm$ 1.01 60.46 $\pm$ 2.06 22.06 $\pm$ 5.58 68.70 $\pm$ 3.63 46.24 $\pm$ 4.83
-BS-KAGCN 75.55 $\pm$ 8.14 73.31 $\pm$ 4.45 71.59 $\pm$ 3.31 75.55 $\pm$ 3.62 57.67 $\pm$ 8.03 70.83 $\pm$ 3.95 50.00 $\pm$ 4.41
-RBF-KAGCN 73.52 $\pm$ 11.6 68.30 $\pm$ 4.53 63.95 $\pm$ 7.80 74.18 $\pm$ 3.64 59.39 $\pm$ 8.03 72.60 $\pm$ 4.61 48.53 $\pm$ 4.03
+GCN 59.95 $\pm$ 8.98 65.87 $\pm$ 6.33 60.10 $\pm$ 4.63 63.10 $\pm$ 3.75 20.56 $\pm$ 3.70 68.40 $\pm$ 4.01 48.56 $\pm$ 3.71
+BS-KAGCN 71.70 $\pm$ 7.54 72.74 $\pm$ 4.16 73.97 $\pm$ 4.01 75.68 $\pm$ 3.49 59.06 $\pm$ 4.48 72.87 $\pm$ 5.12 50.04 $\pm$ 3.99
+RBF-KAGCN 77.86 $\pm$ 9.45 66.78 $\pm$ 5.91 73.07 $\pm$ 2.90 73.67 $\pm$ 3.99 56.83 $\pm$ 8.37 73.13 $\pm$ 4.5 49.84 $\pm$ 3.78
+GAT 57.51 $\pm$ 12.3 71.92 $\pm$ 5.88 60.42 $\pm$ 8.21 59.78 $\pm$ 0.87 18.17 $\pm$ 1.34 67.37 $\pm$ 3.83 48.24 $\pm$ 5.02
+BS-KAGAT 82.66 $\pm$ 4.78 78.27 $\pm$ 2.37 76.46 $\pm$ 3.66 72.75 $\pm$ 2.18 51.33 $\pm$ 12.9 72.00 $\pm$ 3.80 50.09 $\pm$ 4.34
+RBF-KAGAT 75.88 $\pm$ 12.1 73.42 $\pm$ 4.76 72.63 $\pm$ 4.65 71.93 $\pm$ 4.23 49.78 $\pm$ 7.65 71.00 $\pm$ 5.04 48.13 $\pm$ 4.15
+GIN 86.16 $\pm$ 5.96 74.78 $\pm$ 3.82 79.32 $\pm$ 1.63 70.66 $\pm$ 3.60 54.72 $\pm$ 6.55 73.37 $\pm$ 2.90 50.36 $\pm$ 4.10
+BS-KAGIN 85.79 $\pm$ 8.06 76.06 $\pm$ 5.10 80.34 $\pm$ 1.46 73.37 $\pm$ 2.81 44.50 $\pm$ 8.53 73.13 $\pm$ 4.46 50.33 $\pm$ 3.87
+RBF-KAGIN 85.10 $\pm$ 5.60 75.92 $\pm$ 3.49 79.19 $\pm$ 1.81 74.33 $\pm$ 3.52 44.78 $\pm$ 9.58 69.93 $\pm$ 3.84 49.22 $\pm$ 4.41
 Results.
 Table 4 illustrates the average classification accuracies and the corresponding standard deviations of the three models on the different datasets.
-We observe that GIN and both KAGIN  models achieve similar levels of performance on all datasets except ENZYMES.
-In particular, RBF-KAGIN  and BS-KAGIN  both have the best accuracy on $2$ datasets each, and GIN on $1$ dataset.
-Regarding the GCN variants, we observe that the KAN-based variants now have a significant advantage over the linear GCN, confirming that a single KAN layer is more expressive than a simple fully-connexted layer.
-BS-KAGCN  and RBF-KAGCN  obtain the best overall accuracy on $1$ dataset each, while the vanilla GCN performs often significantly poorly.
-These experiments show that, on a graph-level task, which requires significantly more power than node-level tasks, KAN-based models seem to perform better than their MLP-based counterparts.
-While this difference is not significant among GIN-like architectures, it appears that KAN-based GCNs benefit from the higher expressivity in a single layer, and attain performances much closer to their GIN-like counterparts, sometimes even outperforming them.
+We observe that B-splines variants win on $4$ datasets, and the MLP-based GIN leads on $3$, albeit by small margins wrt.
+the standard deviation.
+We notice that, while the MLP-based GCN performs poorly (predictably so, since graph-level tasks often require higher complexity than node-classification ones), the KAN-based GCN perform significantly better.
+This tends to confirm that a single KAN layer has much more representation power than a fully-connected layer.
 5.4 Graph Regression
 Datasets.
-We experiment with two molecular datasets: (1) ZINC-12K [42], and (2) QM9 [43].
+We experiment with two molecular datasets: (1) ZINC-12K (Irwin and Shoichet, 2005), and (2) QM9 (Ramakrishnan et al., 2014).
 ZINC-12K consists of $12,000$ molecules.
@@
 We train each model for $1,000$ epochs by minimizing the mean absolute error (MAE).
-We use the Adam optimizer for model training [39].
+We use the Adam optimizer for model training (Kingma and Ba, 2015).
 We also use early stopping with a patience of $20$ epochs.
@@
 6 Analysis of Size and Time
-5.1 Selected models
+6.1 Setting
 Following the above results, and it is legitimate to wonder whether the performance of the KAN models comes from a larger size than their MLP-based counterparts.
 We show in this section that it is not the case.
-Using the graph-classification models trained for the results provided in 4.2, we display here three relations: the size-to-performance relation, the size-to-training time relation and the training time-to-performance relation.
-Figures 2 and 4 show these relations for datasets NCI1 and IMDB-MULTI respectively.
-We also show in 8 and 8 the numerical values associated to these graphs.
-The rest of the graph classification datasets are provided in App.
-5.2.
-A first thing that we can see from both figures is that, for the same parameter count, B-splines-based KANs are significantly slower than both other architectures.
-RBF-based KANs are much faster, showing that this implementation succeeds in alleviating the computational bottleneck induced by the sequential computation necessary for B-splines.
-Finally, MLPs are the fastest for equal parameter counts.
-While these results were expected, we notice in both the tables and figures that RBF-KAGINmodels end up significantly smaller than their MLP-based counterparts, while maintaining equal or better performance, and about equal training time.
-On the other hand, the B-Splines based method is significantly slower, by sometimes a 5 fold factor.
-This establishes the RBF-based architecture as the best compromise between scalability and performance, although further work could be made in studying the performance of very small KAN-based against MLP-based models.
+Using the graph-classification models trained for the results provided in 5.3, we display here three relations: the size-to-performance relation, the size-to-training time relation and the training time-to-performance relation.
+Note that these models are the ones selected by the hyperparameter search algorithm, thus exhibiting the best validation performance throughout the architecture-optimization process.
+We thus notice an interesting tradeoff, where RBF-based KANs tend to train models with good performance, but smaller size than their MLP-based counterparts, allowing them to train almost as quickly.
+Meanwhile BS-based KANs tend to be slower.
+6.2 Discussion
+Figures 1 shows these relations on dataset NCI1 for the GIN-based models, a setting where the three models have a similar performance.
+We also show in table 6 the associated numerical values for this dataset.
+The rest of the graph classification datasets and models are provided in App.
+C.1.
+We notice that, for the same number of parameters, MLPs are the fastest, followed closely by RBF-based KANs; then BS-based KANs are significantly slower than both others.
+This confirms the fact that the RBF-based representation alleviates greatly the computational bottlenecks of B-splines.
+Nonetheless, we also notice that, despite the hyperparameter search setting allowing them a large size, the performance-driven model selection converged towards smaller sizes for KAN-based models than for MLPs.
+This allows the RBF-based version of KAGNN models to train in time similar to their MLP-based counterpart.
+Nonetheless, the BS-based variant remains significantly slower in almost all settings.
 Figure 1: Parameters count/size/performance comparison, GIN, NCI1
+Table 6: Average time and parameter sizes for selected models, NCI1
+Model Parameter count Accuracy time per epoch (s)
+GCN 316588.40 0.60 $\pm$ 0.04 0.31
+BS-KAGCN 182464.80 0.74 $\pm$ 0.04 0.94
+RBF-KAGCN 117043.10 0.73 $\pm$ 0.03 0.37
+GAT 1566334.00 0.60 $\pm$ 0.08 0.47
+BS-KAGAT 1109714.80 0.76 $\pm$ 0.04 1.42
+RBF-KAGAT 1468335.80 0.73 $\pm$ 0.04 0.53
+GIN 1557454.00 0.79 $\pm$ 0.02 0.39
+BS-KAGIN 532684.50 0.80 $\pm$ 0.01 1.70
+RBF-KAGIN 107075.40 0.79 $\pm$ 0.02 0.37
+7 Discussion and Perspectives
+Prominent application domains.
+KANs, both in our work and in the recent literature, seem to be a promising alternative to MLPs for regression tasks.
+They rely on combinations of continuous functions which could explain their improved performance for continuous outputs.
+Moreover, this property and their empirical performance in regression tasks renders them an ideal candidate for physics-inspired architectures (Karniadakis et al., 2021), which address problems of high importance, from climate forecasting (Kashinath et al., 2021) to atomic simulations (Batzner et al., 2022) and rely heavily on graph-based architectures.
+Explainability matters.
+Recent works on KANs argue that their potential for explainability is higher than that of MLPs with post-hoc methods (e.g.
+Shapley values, gradient-based) (Liu et al., 2025), since closed-form symbolic formulas can be drawn from the former (Liu et al., 2024), providing complete overview of the underlying computations.
+Hence, KAN-based models could prove more friendly to practitioners from application fields, such as biology and physics, who tend to be skeptical of the black-box approach.
+GNNs for fixed graphs inspired from KANs.
+Each KAN architecture consists of a set of neurons.
+Each one of those neurons is connected to other neurons.
+Thus, KANs can be naturally modeled as graphs where nodes represent neurons and edges represent the connections between them.
+Each layer of a KAN is actually a complete bipartite graph, while a KAN has learnable activation functions on the edges of its associated graph.
+There actually exist datasets that contain multiple graphs which share the same graph structure, but differ in the node and edge features.
+Two examples of such datasets are the MNIST and CIFAR10 graph classification datasets (Dwivedi et al., 2023).
+Since the graph structure is fixed (and identical for all graphs), we could associate each edge of the graph with one or more learnable activation functions (shared across graphs), similar to what KANs do.
+This would allow the model to learn different functions for each edge of the graph, leading potentially to high-quality node and graph representations.
+Advantages and future work.
+Finally, we discuss potential advantages that KANs might have over MLPs, and leave their investigation for future work.
+First, their ability to accurately fit smooth functions could prove relevant on datasets where variables interact with some regular patterns.
+Second, their interpretability could be leveraged to provide explanations on learned models, giving insights into the nature of interactions among entities.
+Another study direction could be to extend the KAN architecture to more GNN models (e.g.
+GAT, GraphSAGE, Graph Transformer) to develop a wider KAN-based arsenal for graph tasks.
+Finally, a thorough study of the effect of the different hyperparameters could be pursued, allowing to fully exploit the splines or radial basis functions while retaining small networks.
+8 Conclusion
+In this paper, we have investigated the potential of Kolmogorov-Arnold networks in graph learning tasks.
+Since the KAN architecture is a natural alternative to the MLP, we developed three GNN architectures, KAGCN, KAGAT and KAGIN, respectively analogous to the GCN, GAT and GIN models.
+We then compared those architectures against each other in both node- and graph-level tasks.
+For link prediction, as well as graph and node classification, there does not appear to be a clear winner.
+For graph regression, our results indicate that KAN has an advantage over MLP, with the B-spline variant proving superior in terms of performance.
+Nonetheless, this gain comes at the cost of computational efficiency.
+In this regard, the RBF-based variant offers a lesser performance improvement, but proves much more efficient in terms of training time.
+This paper shows that such KAN-based GNNs are valid alternatives to the traditional MLP-based models.
+We thus believe that these models deserve the attention of the graph machine-learning community, and that future work should be carried out to identify their specific advantages and weaknesses, allowing for an enlightened choice between KANs and MLPs based on the task and context at hand.
+Acknowledgements
+This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation and partner Swedish universities and industry.
+Part of the experiments was enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no.
+2022-06725.
+Part of the experiments was also enabled by the Berzelius resource, which was made possible through application support provided by National Supercomputer Centre at Linköping University.
+Appendix A Further Graph Classification Experiments
+Besides the graph classification experiments that are presented in the main paper, we also experimented with the Peptides-func dataset, a graph classification dataset from the Long Range Graph Benchmark [Dwivedi et al., 2022].
+The dataset consists of $15,535$ graphs where the average number of nodes per graph is equal to $150.94$.
+We evaluated GIN, BS-KAGIN and RBF-KAGIN on this dataset.
+We performed a grid search where for all models we chose the learning rate from $\{0.001,0.0001\}$, the number of hidden layers of MLPs and KANs from $\{1,2,3\}$ and the hidden dimension from $\{64,128\}$.
+We set the number of neighborhood aggregation layers to $5$ for all models.
+For both BS-KAGIN and RBF-KAGIN, grid size was chosen from $\{3,5\}$, and for BS-KAGIN, spline order was chosen from $\{3,5\}$.
+Each experiment was run $4$ times with $4$ different seeds and Table 7 illustrates the average precision achieved by the three models.
+Table 7: Average Precision ($\pm$ standard deviation) of the KAGIN and GIN models on Peptides-func
+Peptides-func
+GIN 55.01 $\pm$ 0.75
+BS-KAGIN
+58.71
+$\pm$ 1.02
+RBF-KAGIN 57.68 $\pm$ 2.05
+We observe that BS-KAGIN and RBF-KAGIN outperform GIN on the Peptides-func dataset.
+Both models offer an absolute improvement of more than $2.5\%$ in average precision over GIN.
+Thus, even in tasks where long-range interactions between nodes need to be captured, for a fixed receptive field, KAN layers can help the models better capture those interactions.
+Appendix B Hyperparameter search ranges
+Table 8: Hyperparameter ranges for node classification (datasets Cora, Cornell, Texas, Wisconsin)
+Model Learning Rate Hidden Layers Hidden Dim. Grid Size Spl. Order Dropout
+GIN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[1,256\right]$ NA NA $\left[0,0.9\right]$
+BS-KAGIN/KAGCN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[1,128\right]$ $\left[1,8\right]$ $\left[1,3\right]$ $\left[0,0.9\right]$
+RBF-KAGIN/KAGCN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[2,128\right]$ $\left[2,32\right]$ NA $\left[0,0.9\right]$
+Table 9: Hyperparameter ranges for node classification (datasets CiteSeer, Ogbn-Arxiv, Actor)
+Model Learning Rate Hidden Layers Hidden Dim. Grid Size Spl. Order Dropout
+GIN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[1,64\right]$ NA NA $\left[0,0.9\right]$
+BS-KAGIN/KAGCN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[2,32\right]$ $\left[1,8\right]$ $\left[1,3\right]$ $\left[0,0.9\right]$
+RBF-KAGIN/KAGCN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[2,32\right]$ $\left[2,4\right]$ NA $\left[0,0.9\right]$
+Table 10: Hyperparameter ranges for link prediction
+Model Learning Rate Hidden Layers Hidden Dim. Grid Size Spl. Order
+GIN $\{1e-2,1e-3,1e-4\}$$\{2,4\}$$\{16,512\}$ NA NA
+[…]
+RBF-KAGIN/KAGCN $\{1e-2,1e-3\}$$\{2,4\}$$\{16,512\}$$\{2,16\}$ NA
+Table 11: Hyperparameter ranges for graph classification
+Model Learning Rate Hidden Layers Hidden Dim. Dropout Grid Size Spl. Order
+GIN $\left[1e-4,1e-2\right]$ $\left[1,4\right]$ $\left[2,512\right]$ $\left[0,0.9\right]$ NA NA
+BS-KAGIN/KAGCN $\left[1e-4,1e-2\right]$ $\left[1,4\right]$ $\left[2,64\right]$ $\left[0,0.9\right]$ $\left[2,32\right]$ $\left[1,4\right]$
+RBF-KAGIN/KAGCN $\left[1e-4,1e-2\right]$ $\left[1,4\right]$ $\left[2,64\right]$ $\left[0,0.9\right]$ $\left[2,16\right]$ NA
+Appendix C Time/Performance/Accuracy analysis
+Figures 2 to 22 and tables 12 to 17 regarding the time/size/accuracy analysis on the graph classification task.
+Table 12: Average time and parameter sizes for selected models, MUTAG
+Model Parameter count Accuracy time per epoch (s)
+GCN 66421.40 0.60 $\pm$ 0.09 0.02
+BS-KAGCN 31991.40 0.72 $\pm$ 0.07 0.03
+RBF-KAGCN 67221.20 0.78 $\pm$ 0.09 0.02
+GAT 1162504.80 0.58 $\pm$ 0.12 0.02
+BS-KAGAT 521626.00 0.83 $\pm$ 0.05 0.04
+RBF-KAGAT 1119860.90 0.76 $\pm$ 0.11 0.02
+GIN 840858.00 0.86 $\pm$ 0.06 0.02
+BS-KAGIN 145455.00 0.86 $\pm$ 0.08 0.05
+RBF-KAGIN 82645.40 0.85 $\pm$ 0.05 0.02
+Table 13: Average time and parameter sizes for selected models, DD
+Model Parameter count Accuracy time per epoch (s)
+GCN 274320.80 0.66 $\pm$ 0.06 0.24
+BS-KAGCN 110847.70 0.73 $\pm$ 0.04 0.93
+RBF-KAGCN 110976.90 0.67 $\pm$ 0.06 0.24
+GAT 224323.60 0.72 $\pm$ 0.06 0.22
+BS-KAGAT 1375974.80 0.78 $\pm$ 0.02 2.36
+RBF-KAGAT 514070.80 0.73 $\pm$ 0.05 0.38
+GIN 281992.60 0.75 $\pm$ 0.04 0.15
+BS-KAGIN 189140.10 0.76 $\pm$ 0.05 2.24
+RBF-KAGIN 62553.50 0.76 $\pm$ 0.03 0.24
+Table 14: Average time and parameter sizes for selected models, PROTEINS_full
+Model Parameter count Accuracy time per epoch (s)
+GCN 14055.70 0.63 $\pm$ 0.04 0.09
+BS-KAGCN 26241.00 0.76 $\pm$ 0.03 0.15
+RBF-KAGCN 77950.00 0.74 $\pm$ 0.04 0.11
+GAT 1217234.00 0.60 $\pm$ 0.01 0.14
+BS-KAGAT 377421.60 0.73 $\pm$ 0.02 0.28
+RBF-KAGAT 406871.00 0.72 $\pm$ 0.04 0.13
+GIN 935245.00 0.71 $\pm$ 0.03 0.10
+BS-KAGIN 206523.00 0.73 $\pm$ 0.03 0.36
+RBF-KAGIN 76445.40 0.74 $\pm$ 0.03 0.14
+Table 15: Average time and parameter sizes for selected models, ENZYMES
+Model Parameter count Accuracy time per epoch (s)
+GCN 230071.20 0.21 $\pm$ 0.04 0.06
+BS-KAGCN 102457.20 0.59 $\pm$ 0.04 0.12
+RBF-KAGCN 105379.80 0.57 $\pm$ 0.08 0.08
+GAT 2193344.40 0.18 $\pm$ 0.01 0.10
+BS-KAGAT 734566.80 0.51 $\pm$ 0.12 0.23
+RBF-KAGAT 292237.60 0.50 $\pm$ 0.07 0.09
+GIN 778435.40 0.55 $\pm$ 0.06 0.07
+BS-KAGIN 199432.90 0.44 $\pm$ 0.08 0.21
+RBF-KAGIN 135783.50 0.45 $\pm$ 0.09 0.10
+Table 16: Average time and parameter sizes for selected models, IMDB-BINARY
+Model Parameter count Accuracy time per epoch (s)
+GCN 106247.80 0.68 $\pm$ 0.04 0.15
+BS-KAGCN 34710.10 0.73 $\pm$ 0.05 0.19
+RBF-KAGCN 40357.50 0.73 $\pm$ 0.04 0.17
+GAT 1482555.60 0.67 $\pm$ 0.04 0.18
+BS-KAGAT 391572.00 0.72 $\pm$ 0.04 0.25
+RBF-KAGAT 308826.00 0.71 $\pm$ 0.05 0.18
+GIN 769820.20 0.73 $\pm$ 0.03 0.16
+BS-KAGIN 144101.00 0.73 $\pm$ 0.04 0.30
+RBF-KAGIN 159608.70 0.70 $\pm$ 0.04 0.20
+Table 17: Average time and parameter sizes for selected models, IMDB-MULTI
+Model Parameter count Accuracy time per epoch (s)
+GCN 83223.60 0.49 $\pm$ 0.04 0.23
+BS-KAGCN 50801.40 0.50 $\pm$ 0.04 0.29
+RBF-KAGCN 25513.00 0.50 $\pm$ 0.04 0.25
+GAT 838081.80 0.48 $\pm$ 0.05 0.25
+BS-KAGAT 129868.00 0.50 $\pm$ 0.04 0.33
+RBF-KAGAT 340916.00 0.48 $\pm$ 0.04 0.27
+GIN 875383.20 0.50 $\pm$ 0.04 0.24
+BS-KAGIN 195207.70 0.50 $\pm$ 0.04 0.38
+RBF-KAGIN 40460.40 0.49 $\pm$ 0.04 0.26
+Figure 2: Parameters count/size/performance comparison, GCN, MUTAG
+Figure 3: Parameters count/size/performance comparison, GAT, MUTAG
+Figure 4: Parameters count/size/performance comparison, GIN, MUTAG
+Figure 5: Parameters count/size/performance comparison, GCN, DD
+Figure 6: Parameters count/size/performance comparison, GAT, DD
+Figure 7: Parameters count/size/performance comparison, GIN, DD
 Figure 8: Parameters count/size/performance comparison, GCN, NCI1
+Figure 9: Parameters count/size/performance comparison, GAT, NCI1
+Figure 10: Parameters count/size/performance comparison, GIN, NCI1
+Figure 11: Parameters count/size/performance comparison, GCN, PROTEINS_full
+Figure 12: Parameters count/size/performance comparison, GAT, PROTEINS_full
+Figure 13: Parameters count/size/performance comparison, GIN, PROTEINS_full
+Figure 14: Parameters count/size/performance comparison, GCN, ENZYMES
+Figure 15: Parameters count/size/performance comparison, GAT, ENZYMES
+Figure 16: Parameters count/size/performance comparison, GIN, ENZYMES
+Figure 17: Parameters count/size/performance comparison, GCN, IMDB-BINARY
+Figure 18: Parameters count/size/performance comparison, GAT, IMDB-BINARY
+Figure 19: Parameters count/size/performance comparison, GIN, IMDB-BINARY
+Figure 20: Parameters count/size/performance comparison, GCN, IMDB-MULTI
+Figure 21: Parameters count/size/performance comparison, GAT, IMDB-MULTI
 Figure 22: Parameters count/size/performance comparison, GIN, IMDB-MULTI
-Figure 4: Parameters count/size/performance comparison, GCN, IMDB-MULTI
-Table 7: Time and parameter sizes for selected models, NCI1
-Model Accuracy Parameter count time per epoch (s)
-BS-KAGIN 0.794 $\pm$ 0.016 572338.4 2.220
-RBF-KAGIN 0.783 $\pm$ 0.017 139168.0 0.358
-GIN 0.783 $\pm$ 0.024 1337369.8 0.359
-BS-KAGCN 0.716 $\pm$ 0.031 90196.5 0.752
-RBF-KAGCN 0.640 $\pm$ 0.075 166960.5 0.356
-GCN 0.604 $\pm$ 0.046 379634.0 0.277
-Table 8: Time and parameter sizes for selected models, IMDB-MULTI
-Model Accuracy Parameter count time per epoch (s)
-BS-KAGIN 0.499 $\pm$ 0.033 296418.8 0.367
-RBF-KAGIN 0.488 $\pm$ 0.043 48705.1 0.169
-GCN 0.503 $\pm$ 0.038 330987.6 0.146
-BS-KAGIN 0.500 $\pm$ 0.042 46052.3 0.216
-RBF-KAGIN 0.485 $\pm$ 0.038 40118.5 0.161
-MLP 0.462 $\pm$ 0.046 111911.4 0.143
-5.2 Theoretical models
+C.1 Training Times - Graph Classification
 We give in Table 18 the training time per epoch for different configurations of the BS-KAGIN, RBF-KAGIN and GIN models on a graph-level task.
@@
 This shows that RBFs do alleviate some of the computational bottlenecks of the B-splines from the original KAN paper.
-Table 16 in Appendix C shows similar observations on a node-level task.
+Table LABEL:tab:timesNodeClassif in Appendix C.2 shows similar observations on a node-level task.
 Table 18: Training time per epoch on NCI1 dataset. $5$ message-passing layers. batch size set to $128$.
@@
 BS-KAGIN 256 3 4 4 10588160 14.572
-6 Discussion
-Prominent application domains.
-KANs, both in our work and in the recent literature, seem to be a promising alternative to MLPs for regression tasks.
-They rely on combinations of continuous functions which could explain their improved performance for continuous outputs.
-Moreover, this property and their empirical performance in regression tasks renders them an ideal candidate for physics-inspired architectures [44], which address problems of high importance, from climate forecasting [45] to atomic simulations [46] and rely heavily on graph-based architectures.
-Explainability matters.
-The potential of explainability for KAN-based models is arguably better compared to that of MLPs with post-hoc methods (e.g.
-Shapley values, gradient-based), since closed-form symbolic formulas can be drawn from the former, providing complete overview of the underlying computations.
-Hence, KAN-based graph models could potentially prove more friendly to practitioners from application fields, such as biology and physics, who tend to be skeptical of the black-box approach.
-GNNs for fixed graphs inspired from KANs.
-Each KAN architecture consists of a set of neurons.
-Each one of those neurons is connected to other neurons.
-Thus, KANs can be naturally modeled as graphs where nodes represent neurons and edges represent the connections between them.
-Each layer of a KAN is actually a complete bipartite graph, while a KAN has learnable activation functions on the edges of its associated graph.
-There actually exist datasets that contain multiple graphs which share the same graph structure, but differ in the node and edge features.
-Two examples of such datasets are the MNIST and CIFAR10 graph classification datasets [47].
-Since the graph structure is fixed (and identical for all graphs), we could associate each edge of the graph with one or more learnable activation functions (shared across graphs), similar to what KANs do.
-This would allow the model to learn different functions for each edge of the graph, leading potentially to high-quality node and graph representations.
-Advantages and future work.
-Finally, we discuss potential advantages that KANs might have over MLPs, and leave their investigation for future work.
-First, their ability to accurately fit smooth functions could prove relevant on datasets where variables interact with some regular patterns.
-Second, their interpretability could be leveraged to provide explanations on learned models, giving insights into the nature of interactions among entities.
-Finally, a thorough study of the effect of the different hyperparameters could be pursued, allowing to fully exploit the splines or radial basis functions while retaining small networks.
-7 Conclusion
-In this paper, we have investigated the potential of Kolmogorov-Arnold networks in graph learning tasks.
-Since the KAN architecture is a natural alternative to the MLP, we developed two GNN architectures, KAGCN and KAGIN, respectively analogous to the GCN and GIN models.
-We then compared those architectures against each other in both node- and graph-level tasks.
-For graph and node classification, there does not appear to be a clear winner.
-For graph regression, our results indicate that KAN has an advantage over MLP, with the B-spline variant proving superior in terms of performance.
-Nonetheless, this gain comes at the cost of computational efficiency.
-In this regard, the RBF-based vairant offers a slightly lesser performance improvement, but proves much more efficient in terms of training time.
-This paper shows that such KAN-based GNNs are valid alternatives to the traditional MLP-based models.
-We thus believe that these models deserve the attention of the graph machine-learning community, and that future work should be carried out to identify their specific advantages and weaknesses, allowing for an enlightened choice between KANs and MLPs based on the task and context at hand.
-Appendix A Further Graph Classification Experiments
-Besides the graph classification experiments that are presented in the main paper, we also experimented with the Peptides-func dataset, a graph classification dataset from the Long Range Graph Benchmark [48].
-The dataset consists of $15,535$ graphs where the average number of nodes per graph is equal to $150.94$.
-We evaluated GIN, BS-KAGIN and RBF-KAGIN on this dataset.
-We performed a grid search where for all models we chose the learning rate from $\{0.001,0.0001\}$, the number of hidden layers of MLPs and KANs from $\{1,2,3\}$ and the hidden dimension from $\{64,128\}$.
-We set the number of neighborhood aggregation layers to $5$ for all models.
-For both BS-KAGIN and RBF-KAGIN, grid size was chosen from $\{3,5\}$, and for BS-KAGIN, spline order was chosen from $\{3,5\}$.
-Each experiment was run $4$ times with $4$ different seeds and Table 10 illustrates the average precision achieved by the three models.
-Table 10: Average Precision ($\pm$ standard deviation) of the KAGIN and GIN models on the Peptides-func dataset.
-Peptides-func
-GIN 55.01 $\pm$ 0.75
-BS-KAGIN
-58.71
-$\pm$ 1.02
-RBF-KAGIN 57.68 $\pm$ 2.05
-We observe that BS-KAGIN and RBF-KAGIN outperform GIN on the Peptides-func dataset.
-Both models offer an absolute improvement of more than $2.5\%$ in average precision over GIN.
-Thus, even in tasks where long-range interactions between nodes need to be captured, for a fixed receptive field, KAN layers can help the models better capture those interactions.
-Appendix B Time/Performance/Accuracy analysis
-Figures 6 to 14 and tables 11 to 15 regarding the time/size/accuracy analysis on the graph classification task.
-Table 11: Time and parameter sizes for selected models, IMDB-BINARY
-Model Accuracy Parameter count time per epoch (s)
-BS-KAGIN 0.734 $\pm$ 0.038 148236.3 0.225
-RBF-KAGIN 0.720 $\pm$ 0.036 168383.6 0.129
-GIN 0.731 $\pm$ 0.038 1172520.5 0.110
-BS-KAGCN 0.708 $\pm$ 0.037 39053.5 0.150
-RBF-KAGCN 0.726 $\pm$ 0.044 37322.5 0.109
-GCN 0.687 $\pm$ 0.034 118826.3 0.099
-Table 12: Time and parameter sizes for selected models, ENZYMES
-Model Accuracy Parameter count time per epoch (s)
-BS-KAGIN 0.427 $\pm$ 0.085 450136.6 0.336
-RBF-KAGIN 0.412 $\pm$ 0.079 115346.8 0.061
-GIN 0.533 $\pm$ 0.117 1623033.2 0.048
-BS-KAGCN 0.577 $\pm$ 0.076 83636.3 0.103
-RBF-KAGCN 0.594 $\pm$ 0.053 125854.4 0.050
-GCN 0.221 $\pm$ 0.053 280040.0 0.039
-Table 13: Time and parameter sizes for selected models, DD
-Model Accuracy Parameter count time per epoch (s)
-BS-KAGIN 0.754 $\pm$ 0.032 169417.4 1.206
-RBF-KAGIN 0.757 $\pm$ 0.029 147005.1 0.236
-GIN 0.748 $\pm$ 0.044 168957.4 0.144
-BS-KAGCN 0.733 $\pm$ 0.042 110278.6 0.596
-RBF-KAGCN 0.683 $\pm$ 0.043 110899.2 0.185
-GCN 0.632 $\pm$ 0.036 342023.2 0.225
-Table 14: Time and parameter sizes for selected models, MUTAG
-Model Accuracy Parameter count time per epoch (s)
-BS-KAGIN 0.851 $\pm$ 0.058 236235.8 0.056
-RBF-KAGIN 0.855 $\pm$ 0.051 62530.0 0.020
-GIN 0.853 $\pm$ 0.069 848213.0 0.014
-BS-KAGCN 0.755 $\pm$ 0.077 19496.1 0.022
-RBF-KAGCN 0.735 $\pm$ 0.110 47769.4 0.017
-GCN 0.563 $\pm$ 0.143 72072.2 0.012
-Table 15: Time and parameter sizes for selected models, PROTEINS_full
-Model Accuracy Parameter count time per epoch (s)
-BS-KAGIN 0.741 $\pm$ 0.034 321367.7 0.472
-RBF-KAGIN 0.733 $\pm$ 0.034 51498.8 0.096
-MLP 0.710 $\pm$ 0.027 942781.0 0.073
-BS-KAGCN 0.756 $\pm$ 0.034 37688.7 0.133
-RBF-KAGCN 0.742 $\pm$ 0.035 48950.4 0.078
-GCN 0.605 $\pm$ 0.020 8199.8 0.061
-Figure 5: Parameters count/size/performance comparison, GIN, IMDB-BINARY
-Figure 6: Parameters count/size/performance comparison, GCN, IMDB-BINARY
-Figure 7: Parameters count/size/performance comparison, GIN, ENZYMES
-Figure 8: Parameters count/size/performance comparison, GCN, ENZYMES
-Figure 9: Parameters count/size/performance comparison, GIN, DD
-Figure 10: Parameters count/size/performance comparison, GCN, DD
-Figure 11: Parameters count/size/performance comparison, GIN, MUTAG
-Figure 12: Parameters count/size/performance comparison, GCN, MUTAG
-Figure 13: Parameters count/size/performance comparison, GIN, PROTEINS_full
-Figure 14: Parameters count/size/performance comparison, GCN, PROTEINS_full
-Appendix C Training Times - Node Classification
+C.2 Training Times - Node Classification
 Table 19: Training time per epoch on the node-classification task (Cora dataset, 3 MP-layers).